### v1

# An Effective Version of the $p$-Curvature Conjecture for Order One Differential Equations

An Effective Version of the $p$-Curvature Conjecture
for Order One Differential Equations
Florian Fьrnsinn and Lucas Pannier
Аннотация
We develop an effective version of Kronecker’s Theorem on the splitting of polynomials, based on asymptotic arguments proposed by the Chudnovsky brothers, coming from Hermite-Padй approximation. In conjunction with Honda’s proof of the $p$-curvature conjecture for order one equations with polynomial coefficients we use this to deduce an effective version of the Grothendieck $p$-curvature conjecture for order one equations. More precisely, we bound the number of primes for which the $p$-curvature of a given differential equation has to vanish in terms of the height and the degree of the coefficients, in order to conclude it has a non-zero algebraic solution. Using this approach, we describe an algorithm that decides algebraicity of solutions of differential equation of order one using $p$-curvatures, and report on an implementation in SageMath.
footnotetext: MSC 2020 classification: 34M15 (primary); 11Y05, 11Y16, 68W30 (secondary)
1 Introduction
A power series $f(x)\in\mathbb{Q}[\![x]\!]$ is called algebraic if there exists a non-zero polynomial $P(x,y)\in\mathbb{Q}[x,y]$, such that $P(x,f(x))=0$. It is a well-known fact, already known to Abel, that every algebraic power series satisfies a linear differential equation
$a_{n}(x)y^{(n)}(x)+a_{n-1}(x)y^{(n-1)}(x)+\ldots+a_{1}(x)y^{\prime}(x)+a_{0}(x)y(x)=0$ (1)
with polynomial coefficients $a_{i}(x)\in\mathbb{Q}[x]$ for $i=0,\ldots,n.$
Conversely, deciding whether the solutions of (1) are algebraic is an old problem dating back at least to Fuchs, and Liouville. More precisely, given (1), one might ask, if
  • (A)
    all solutions are algebraic,
  • (E)
    there exists at least one non-zero solution that is algebraic, or
  • (P)
    a particular solution, given for instance by enough coefficients of its power series expansion such that it is uniquely determined, is algebraic.
Notably, in his paper from 1980 that popularized the study of D-finite series, i.e., power series that satisfiy a non-zero differential equation of the form (1), Stanley asked for an ‘‘algorithm suitable for computer implementation’’ for problem (P) [Sta80].
Problem (A) was solved algorithmically in a seminal paper by Singer [Sin79], relying on an algorithmic answer of Risch [Ris70] in the case of first order differential equations with algebraic power series coefficients instead of polynomials – this instance of the problem is known as Abel’s problem – and on work conducted by Painlevй and Boulanger [Pai87, Bou97]. However, this answer is not fully satisfactory, as the complexity of the algorithm involves exponential bounds, making it in general not suitable for implementation. Recent progress on Stanley’s question about problem (P) was made by Bostan, Salvy and Singer [BSS25], giving (semi-) algorithms that rely on the minimization of differential operators annihilating a given D-finite series.
A different – arithmetic – approach to problem (A) was proposed by Grothendieck, in his $p$-curvature conjecture [Kat72]. Given (1), we can pass to a matrix differential equation $Y^{\prime}(x)=A(x)Y(x)$ of order $1$, with $A(x)\in\mathbb{Q}(x)^{n\times n}$ being the companion matrix of (1). Let $p$ be a prime number. The $p$-curvature of the equation is defined as the $\mathbb{F}_{p}(x)$-linear map $\psi_{A}:\mathbb{F}_{p}(x)^{n}\to\mathbb{F}_{p}(x)^{n},Y(x)\mapsto(\partial-A_{p})^{p}Y(X)$, where $A_{p}(x)\in\mathbb{F}_{p}(x)^{n\times n}$ denotes the reduction of $A(x)$ modulo $p$. The $p$-curvature conjecture then states that (1) admits a basis of algebraic solutions if and only if its $p$-curvature vanishes for almost all, i.e., all but finitely many, prime numbers $p$. Cartier’s Lemma (as attributed by Katz [Kat72]), asserts that the vanishing of the $p$-curvature is equivalent to the existence of a full basis of (algebraic) solutions of the reduction of the differential equation modulo $p$ in $\mathbb{F}_{p}[\![x]\!]$. Thus, the $p$-curvature conjecture can be seen as a local-global principle about the existence of algebraic solutions of a linear differential equation.
While in general wide open, the $p$-curvature conjecture is solved in many cases. Most notably, for ‘‘equations coming from geometry’’, i.e., suitable factors of Picard-Fuchs differential operators, a proof was given by Katz [Kat72], and for hypergeometric differential equations an elementary proof is a by-product of the classification of algebraic hypergeometric functions [BH89, FY24]. For first order equations
$y^{\prime}(x)=u(x)y(x)$ (2)
with $u(x)\in\mathbb{Q}(x)$ the three problems (A), (E) and (P) are equivalent, as there is only a one dimensional solution space of the equation. Honda [Hon81] provided an elementary proof of the $p$-curvature conjecture in this case, by showing that the problem is equivalent to a number theoretic result by Kronecker [Kro80], which nowadays is often seen as a consequence of Chebotarev’s Density Theorem [Чеб23a, Чеб23, Tsc26].
Theorem 1.1 (Kronecker).
Let $R(x)\in\mathbb{Q}[x]$ be an irreducible polynomial. If for almost all prime numbers $p$, the reduction of $R(x)$ modulo $p$ has a root in $\mathbb{F}_{p}$, then $R(x)$ has a root in $\mathbb{Q}$, hence $R(x)$ is linear.
For first order equations with algebraic power series coefficients, i.e., for the equation $y^{\prime}(x)=u(x)y(x)$ with $u(x)\in\overline{\mathbb{Q}}[\![x]\!]\cap\overline{\mathbb{Q}(x)}$, a proof of the $p$-curvature conjecture was given by the Chudnovsky brothers [CC85]. They use Hermite-Padй approximation to obtain their result. As a motivating example, they apply their methods to provide a new proof of Kronecker’s Theorem, not relying on Chebotarev’s Density Theorem, and in doing so giving also a new proof of Honda’s Theorem.
The $p$-curvature conjecture, as stated here, does not provide a way to decide whether (1) has a basis of algebraic solutions, as it translates the statement to a statement about almost all, in particular, a statement about infinitely many, prime numbers. In 1982 Katz extended the $p$-curvature conjecture to a conjecture about the Lie algebra of the differential Galois group of a differential equation [Kat82]. From his work presented in [Kat82, §9] it follows by a noetherianity argument that - in the cases in which his conjecture is true, among them order one equations with rational function coefficients - there exists a finite set of prime numbers, such that the vanishing of the $p$-curvatures modulo these primes implies algebraicity of the solutions. However, his argument is non-constructive. The purpose of this article is to explicitly construct a set of primes, for which it suffices to check the vanishing of the $p$-curvatures of a first order equation (2) – the cardinality of the set depending on a suitable measure of the ‘‘size’’ of $u(x)$ – to conclude that it has a non-zero algebraic solution. Our first main result is the following effective version of Kronecker’s Theorem.
Theorem 1.2.
Let
$R(w)=r_{n}w^{n}+\ldots+r_{1}w+r_{0}\in\mathbb{Z}[w]$
be a polynomial with leading coefficient $\Delta\coloneqq r_{n}>0$, and suppose that the maximal modulus of its complex roots is bounded by $B\in\mathbb{R}$. Let $\delta(\Delta)\coloneqq\prod_{p|\Delta}p^{1/(p-1)}$ and set $M\coloneqq\left\lceil 2.826\cdot\Delta^{3}\cdot\delta(\Delta)^{3}\right\rceil$ and $N\coloneqq\lceil 6.076\cdot BM\rceil$. Then $R(w)$ splits into linear factors in $\mathbb{Q}[w]$ if and only if its reduction modulo any prime number $p$ not dividing $\Delta$ and less than $\sigma\coloneqq(2M+1)N+2M$ splits into linear factors in $\mathbb{F}_{p}[w]$.
Our proof of this result makes the Chudnovskys’ proof of Kronecker’s result more explicit. The method of Hermite-Padй approximation inherently is effective in some sense, and in the conclusion of their article [CC85], the Chudnovsky brothers also claim that their results are effective. However most of their computations are sketched without many details, preventing us from directly deriving explicit bounds. Our contribution is to confirm their claim and to work out concrete bounds.
With this we obtain as a corollary the following method of deciding algebraicity of solutions of order one differential equations.
Theorem 1.3.
Let $u(x)=c\cdot\frac{a(x)}{b(x)}\in\mathbb{Q}(x)$ be a rational function with $a(x),b(x)\in\mathbb{Z}[x]$ primitive [^1] Recall that a polynomial $a(x)=a_{n}x^{n}+\ldots+a_{1}x+a_{0}\in\mathbb{Z}[x]$ is called primitive, if $\gcd\{a_{0},\ldots,a_{n}\}=1$., and $c\in\mathbb{Q}$. Assume that the coefficients of $a(x)$ and $b(x)$ are bounded in absolute value by $H$. Let
$R(w)\coloneqq\operatorname{res}_{x}(b(x),a(x)-w\cdot b^{\prime}(x))=r_{n}w^{n}+\ldots+r_{1}w+r_{0}$
and let $\Delta\coloneqq|\operatorname{res}(b(x),b^{\prime}(x))|=|r_{n}|$. With the notation as in Theorem 1.2, the equation $y^{\prime}(x)=u(x)y(x)$ has a non-zero algebraic solution if and only if its $p$-curvatures vanish for all primes not dividing $\Delta$ and smaller than $\sigma$.
The computational complexity of checking that sufficiently many $p$-curvatures vanish is $\tilde{O}(\Delta^{6}B)=\tilde{O}(H^{12n-6}n^{12n}3^{-3n})$ where the notation $\tilde{O}$ hides factors that are polynomial in $n$ and logarithmic in $H$.
We also explain how the problem of deciding algebraicity of solutions of order one differential equations with algebraic coefficients reduces to the case of rational coefficients as in the theorem above.
Moreover, we treat algorithmic aspects of the result and report on an implementation. We rely on a fast algorithm by Bostan and Schost [BS09] to quickly compute $p$-curvatures of order one equations. The algorithm presented in this text does not outperform other known algorithms for certifying that a differential equation of order one has algebraic solutions, however in generic cases the algorithm detects quickly the presence of a transcendental solution, especially when the degree and size of the coefficients explode. We expect that any significant improvement in our approach for certifying the algebraicity of solutions would come from new theoretical results rather than algorithmic optimization. Although theoretically the algorithm could treat the case of differential equations with algebraic non-rational coefficients, we did not optimize it to treat this case efficiently.
Our algorithm provides an instance of deciding algebraicity of D-finite functions using exclusively an arithmetic criterion. One could hope for results, similar in spirit, for differential equations of order one with algebraic coefficients, and for higher order equations in which the $p$-curvature conjecture is proven. The computation of $p$-curvatures for arbitrary order differential equations and deciding their nullity is possible and algorithms exist to perform these computations efficiently [BCS15, BCS16].
Structure of the Paper.
In Section 2 we revisit Honda’s proof of the $p$-curvature conjecture for order one differential equations by investigating its equivalence with Kronecker’s Theorem in view of effective aspects. Relying on Theorem 1.2, we give a proof of Theorem 1.3, with the exception of the stated computational complexity. In Section 3 we discuss different effective approaches to the $p$-curvature conjecture for order one equations. The proof of Theorem 1.2 takes up the entirety of Section 4, in which we adapt the Chudnovskys’ proof of Kronecker’s Theorem. Afterwards, in Section 5, we describe an algorithm that decides the algebraicity of the solution of an order one differential equation using the computation of $p$-curvatures. Its complexity estimate, worked out in Proposition 5.15 finishes the proof of Theorem 1.3. Finally, in the last section, Section 6, we discuss our implementation of this algorithm in SageMath.
Acknowledgments.
We are indebted to Alin Bostan, for his continuous support, for lots of insightful discussions and for many useful references. We thank Markus Reibnegger for his interest in the work, for pointing out crucial references, and for fruitful discussions.
The first-named author was funded by a DOC Fellowship (27150) of the Austrian Academy of Sciences (ЦAW) at the University of Vienna. The second-named author was funded by a CNRS MITI grant. Both authors were supported by the French–Austrian project EAGLES (ANR-22-CE91-0007 & FWF grant 10.55776/I6130).
2 Reformulations of the Problem
This section contains no new results. Its purpose is to describe the equivalence between the $p$-curvature conjecture for the order one equation (2) and Kronecker’s Theorem 1.1. The results can be found in one form or another in the literature [Hon81, CC85, Put96, BCR24].
We consider the equation (2) with $u(x)\in\mathbb{Q}(x)$. A nonzero solution $y(x)$ has the form $y(x)=\exp(\int u(x)\mathrm{d}x)$. This expression combines two operations that do not preserve rationality or algebraicity in different ways. A primitive of a rational function remains a rational function if and only if all its residues are zero, whereas a fraction $\frac{\alpha}{x-\beta}$ has a logarithm as a primitive. Contrarily, the exponential acts in such a way that for a nonzero algebraic function $f(x)$, the function $\exp(f(x))$ is transcendental. This can be seen as a particular case of a conjecture by Schanuel about the transcendence degree of a set of power series and their exponentials, that was proven by Ax [Ax71].
Hence, for a function $y(x)=\exp(\int u(x)\mathrm{d}x)$ to be algebraic, $u(x)$ cannot have a polynomial part or poles of order more than one, so that $y(x)$ factors into a product of $(x-\beta)^{\alpha}$ where $\beta$ is a pole of $u(x)$, and $\alpha$ is the residue at $\beta$. For such $u(x)$, the function $y(x)$ is algebraic if and only if the residues of $u(x)$ are rational numbers.
A convenient way to compute the residues of the rational function $u(x)$ is provided by Rothstein-Trager resultants [Rot76, Rot77, Tra76]. We state their result here in a simplified, but for our purposes sufficient, form.
Theorem 2.1 (Rothstein, Trager).
Let $u(x)=a(x)/b(x)$ be a rational function with only simple poles and $\deg a(x)<\deg b(x)$. Then its residues are precisely the roots of the Rothstein-Trager resultant
$R(w)\coloneqq\operatorname{res}_{x}(b(x),a(x)-w\cdot b^{\prime}(x)).$
We investigate complexity estimates for Rothstein-Trager resultants in Section 5.1. Altogether we get the following equivalent characterizations.
Proposition 2.2.
The following statements for the differential equation (2) are equivalent.
  1. (1)
    All solutions of (2) are algebraic.
  2. (2)
    The coefficient $u(x)$ can be written as
    $u(x)=\sum\limits_{i=1}^{n}\frac{\alpha_{i}}{x-\beta_{i}}$ (3)
    with distinct algebraic poles $\beta_{i}\in\overline{\mathbb{Q}}$ and rational residues $\alpha_{i}\in\mathbb{Q}$.
  3. (3)
    We have $\deg a(x)<\deg b(x)$, all poles of $u(x)$ are simple and its Rothstein-Trager resultant factors into linear factors in $\mathbb{Q}[w]$.
The conditions on the degree and the poles of $u(x)$ in the third item are equivalent to the differential equation (2) being Fuchsian.
Similar considerations hold in characteristic $p$. Moreover, by Cartier’s Lemma, the existence of a non-zero solution of an order one differential equation $y^{\prime}(x)=u(x)y(x)$ is equivalent to the vanishing of its $p$-curvature [^2] The existence of a non-zero polynomial solution in $\mathbb{F}_{p}[x]$, a non-zero algebraic solution in $\mathbb{F}_{p}[\![x]\!]$ or any non-zero solution in $\mathbb{F}_{p}[\![x]\!]$ are equivalent [Hon81, Lem. 1].. In the particular case of order one equations, the $p$-curvature is given by a one-dimensional matrix, whose entry we also call, by abuse of notation, the $p$-curvature. There is an explicit formula for this $p$-curvature in terms of the coefficient $u(x)$, that already implicitly appears in Jacobson’s work [Jac37]: it is given by $u(x)^{p}+u^{(p-1)}(x)$.
Putting all these observations together, we infer the following results characterizing the existence of (algebraic) solutions in characteristic $p$.
Proposition 2.3.
Let $p$ be a prime number. The following statements are equivalent.
  1. (1)
    The $p$-curvature $u(x)^{p}+u^{(p-1)}(x)$ of (2) vanishes.
  2. (2)
    The reduction of (2) has an algebraic solution in $\mathbb{F}_{p}[\![x]\!]$.
  3. (3)
    The coefficient $u(x)$ can be written as
    $u(x)=\frac{a(x)}{b(x)}=\sum_{i=1}^{n}\frac{\alpha_{i}}{x-\beta_{i}},$
    with $\beta_{i}\in\overline{\mathbb{F}}_{p}$ and $\alpha_{i}\in\mathbb{F}_{p}$.
  4. (4)
    We have $\deg a(x)<\deg b(x)$, all poles of $u(x)$ are simple and its Rothstein-Trager resultant factors into linear factors in $\mathbb{F}_{p}[w]$.
Honda’s proof of the $p$-curvature conjecture for (2) essentially consists in recognizing that Kronecker’s Theorem implies that if (4) in Proposition 2.3 holds for almost all prime numbers $p$, then (3) in Proposition 2.2 holds as well. Our following considerations throughout this article are focused on the question, whether (4) in Proposition 2.3 for finitely many prime numbers suffices to conclude (3) in Proposition 2.2.
Note that Honda actually proved more, namely that the $p$-curvature conjecture for order one equations is equivalent to Kronecker’s Theorem. We refer the reader for the other implication to Honda’s original article [Hon81].
Let us next investigate which prime numbers are exceptional in the sense that there is no solution of the reduction of the differential equation $y^{\prime}(x)=u(x)y(x)$ with $u(x)=a(x)/b(x)$ in $\mathbb{F}_{p}[\![x]\!],$ despite the existence of a solution in characteristic 0. In other words we are investigating for which prime numbers assertion (4) in Proposition 2.3 might be false, although (3) in Proposition 2.2 holds. Trivially, if a polynomial splits in $\mathbb{Z}[x]$ into linear factors, it splits into linear factors in $\mathbb{F}_{p}[x]$ not just for almost all prime numbers, but for all prime numbers. However, it might happen that the degree of $b(x)$ drops, when reduced modulo $p$, or that two distinct poles of $u(x)$ of order $1$ collapse to a pole of higher order, when reduced modulo $p$. Luckily, both of these exceptions can be easily read off the leading coefficient (up to sign) $\Delta=|\operatorname{res}(b(x),-b^{\prime}(x))|$ of the Rothstein-Trager resultant, as already noticed by van der Put, c.f. [Put96, Prop. 3.2(2)].
Proposition 2.4.
Assume that $y^{\prime}(x)=u(x)y(x)$ with $u(x)=a(x)/b(x)$ has an algebraic solution. If $p\not\hskip 2.5pt\mid\Delta$ then the $p$-curvature of the equation vanishes.
The following example illustrates both types of exceptional primes.
Example 2.5.
Consider the rational function
$u(x)=\frac{a(x)}{b(x)}=\frac{x+2}{2x^{2}+x-1}=\frac{5}{6}\cdot\frac{1}{x-\frac{1}{2}}-\frac{1}{3}\cdot\frac{1}{x+1}$
and the associated differential equation $y^{\prime}(x)=u(x)y(x)$. It has the algebraic solution $y(x)=\frac{\left(2x-1\right)^{{5}/{6}}}{\left(x+1\right)^{{1}/{3}}}$, which cannot be reduced modulo $2$ and $3$, but modulo all other prime numbers. The resultant $\operatorname{res}(b(x),-b^{\prime}(x))$ is equal to 18-18. For $p=2$ the denominator of $u(x)$ reduces to a degree one polynomial. For $p=3$ the two poles $x=\frac{1}{2}$ and x=1x=-1 collapse to a single pole of order $2$.
We are now ready to prove Theorem 1.3, assuming Theorem 1.2.
Proof of Theorem 1.3.
First, we can reduce without loss of generality to the case $c=1$. Indeed, if $y(x)$ satisfies $y^{\prime}(x)=u(x)y(x)$, then $(y(x)^{c})^{\prime}=c\cdot u(x)y(x)^{c}$ and for a given $c\in\mathbb{Q}$, the function $y(x)$ is algebraic if and only if $y(x)^{c}$ is algebraic.
By Proposition 2.2 it suffices to check that $\deg b(x)>\deg a(x)$, that $b(x)$ only has simple poles, and that $R(w)=\operatorname{res}_{x}(b(x),a(x)-w\cdot b^{\prime}(x)$ factors in linear factors in $\mathbb{Q}[w]$. If one of the first two conditions is not met, it is also not met when reducing $u(x)$ modulo almost all primes, so by Proposition 2.3 almost all $p$-curvatures will not vanish.
By Theorem 1.2, $R(w)$ factors into linear factors in $\mathbb{Q}[w]$ if and only if its reduction also splits completely in $\mathbb{F}_{p}[w]$ for all prime numbers $p$ up to $\sigma$, not dividing $\Delta$. By Proposition 2.3 this in turn is equivalent to the vanishing of the $p$-curvatures for all prime numbers not dividing $\Delta$ and smaller than $\sigma$.
The complexity estimates follow from Proposition 5.15. ∎
Finally we note that our arguments can easily be extended to equations $y^{\prime}(x)=u(x)y(x)$ with $u(x)\in\overline{\mathbb{Q}}[x]$ by using the following well-known result, that is a direct consequence of Vieta’s formulas.
Proposition 2.6.
Let $K$ be a number field, let $R(w)\in K[w]$ with leading coefficient $\Delta\in K\backslash\{0\}$. If $R(w)$ splits completely over $K$ with all its roots in $\mathbb{Q}$, then $\frac{1}{\Delta}R(w)$ is a polynomial in $\mathbb{Q}[w]$.
3 Comparison to other Methods
3.1 Finding Rational Roots
Once we have computed the Rothstein-Trager resultant $R(w)\in\mathbb{Q}[w]$ associated to $u(x)$, checking the rationality of the residues (and thus solving the problem of deciding algebraicity of the solutions of (2)) amounts to finding the rational roots of $R(w)$. Then the residues of the rational function $u(x)$ are all rational if and only if its Rothstein-Trager resultant $R(w)$ splits completely over $\mathbb{Q}$.
A naïve approach to finding the rational roots of a polynomial is to factor it completely over $\mathbb{Q}$ and consider the factors of degree one. However this is more than what we are looking for and specific more efficient algorithms to find the rational roots of a polynomial exist. We refer the reader to [vzGG13, §14-16] and [Bos+17, §21] for details about such algorithms and an analysis of their complexity.
We obtain Algorithm 1.
Algorithm 1 Deciding algebraicity by finding rational roots
[…]
2:Find rational roots of $R(w)$;
3:if there are $\deg(R(w))$ rational roots then return Algebraic;
4:else return Transcendental;
Remark 3.1.
Starting from a rational function in $u(x)\in\overline{\mathbb{Q}}[x]$, deciding the algebraicity of solutions of $y^{\prime}(x)=u(x)y(x)$ reduces to the same problem. Following Proposition 2.6, once we have computed the Rothstein-Trager resultant, we can divide by its leading coefficient and if the obtained polynomial has non-rational coefficients we can conclude on the transcendence of the solutions of the input differential equation. Else we continue with the polynomial in $\mathbb{Q}[w]$ as described in this section.
3.2 The Least Prime That Does Not Split in a Number Field
The matter of bounding the least rational prime number that does not split completely in a given (Galois) number field has been extensively studied in the literature. It is a special case of the more general problem of finding an upper bound on the least prime that has a prescribed splitting pattern in a number field [LMO79, AK19, KW22], in accordance with the Chebotarev Density Theorem. More precisely, given a Galois field extension $K/\mathbb{Q}$ of degree $d$ and discriminant $D$, there are several asymptotic estimates on the size of the smallest rational prime that does not split completely in $K$. An explicit statement was proven by Vaaler and Voloch [VV00, Thm. 1].
Theorem 3.2 (Vaaler, Voloch).
Let $K/\mathbb{Q}$ be a number field of degree $d$ and discriminant $D$. If $\exp(\max\{105,25(\log(d))^{2}\})\leq 8D^{\frac{1}{2(d-1)}}$ then there exists a prime $p$, such that $p$ does not split completely in $K$ and $p\leq 26d^{2}D^{\frac{1}{2(d-1)}}$.
Sharper, asymptotic bounds were proven, both under the assumption of the Generalized Riemann Hypothesis by Murty [Mur94], and unconditionally [Mur94, Li12].
Note that for a prime $p$ splitting completely in the splitting field of a polynomial $R(w)$ is equivalent to the reduction of $R(w)$ modulo $p$ splitting into linear factors in $\mathbb{F}_{p}[w]$. Thus, this gives another approach to finding an effective version of Kronecker’s Theorem. There seem to exist obstacles, however.
First, most results in this direction are of asymptotic nature, making it hard to convert them into an explicit bound. The exception is Vaaler and Voloch’s statement given above. However, they have an assumption on the minimal size of $D$ that exceeds the computational capacity of modern computers. Li claims [Li12, p. 1062] that this assumption is artificial and better results could be expected for small $D$.
Second, one would need to estimate, given a polynomial $R(w)$, the discriminant and the degree of a potential splitting field of $R(w)$. Given that we cannot assume $R(w)$ to be irreducible, the discriminant of the polynomial and the discriminant of its splitting field are not related in an obvious way.
While both of these problems seem to be manageable in principle, we are not aware of any algorithm for solving our problem based on this approach.
3.3 Indicial Equations
As discussed in Section 2, the problem of deciding algebraicity of the solutions of (2) can be solved by investigating its singularities, and, in particular by deciding whether the residues of the rational function coefficient $u(x)$ are rational. Working directly in the framework of differential equations instead, we have the following addition to Proposition 2.2.
Proposition 3.3.
The following statements for the differential equation (2) are equivalent.
  1. (1)
    All solutions of (2) are algebraic.
  2. (4)
    All singularities of (2) are regular singular and its local exponent at each of these singularities is rational.
Доказательство.
The singularities of (2) are precisely the poles of $u(x)$, and possibly $\infty$. As the order of the differential equation is $1$, a singularity at $x_{0}$ is regular, if and only if the valuation of $u(x)$ at $x_{0}$ is at least 1-1, i.e., if the pole of $u(x)$ is of order at most $1$. The local exponent at $x_{0}$ is precisely given by the residue of $u(x)$ at $x_{0}$. Investigating the singularity at $\infty$ amounts to investigate the behavior of $x^{2}y^{\prime}(x)=u(\frac{1}{x})y(x)$ at 0. So the singularity is present, if and only if $\deg a(x)\geq\deg b(x)-1$ and it is regular if and only if $\deg a(x)=\deg b(x)-1$. The local exponent at $\infty$ automatically is rational if $u(x)$ has rational coefficients. So we have shown that (4) is equivalent to assertion (2) of Proposition 2.2. ∎
To use this equivalent criterion to decide algebraicity of solutions, one can factor $b(x)$ to determine the singularities of the equation and then compute the local exponents by computing the indicial polynomial at the zeroes of $b(x)$. This polynomial is of degree one with coefficients in the splitting field of $b(x)$. Its root is rational, if and only if it is a $\overline{\mathbb{Q}}$-multiple of a polynomial in $\mathbb{Q}[x]$.
This approach leads to the following algorithm.
Algorithm 2 Deciding algebraicity with indicial equations
[…]
3:  Compute the indicial equation at $x=z$;
4:  if the root of the indicial equation is not rational then return Transcendental;
5:return Algebraic;
4 An Effective Version of Kronecker’s Theorem
This section is devoted to the proof of Theorem 1.2. We outline the steps here. In [CC85] D.V. and G.V. Chudnovsky provide an elementary proof of Kronecker’s Theorem (and with it, according to the discussion in Section 2, of Honda’s proof of the $p$-curvature conjecture for order one equations) using Hermite-Padй approximation.
One main ingredient is the well-known fact that the function $x^{\alpha}$, for example defined, setting $z\coloneqq x-1$, by the power series
$x^{\alpha}=\sum\limits_{n\geq 0}\binom{\alpha}{n}(x-1)^{n}=\sum\limits_{n\geq 0}\binom{\alpha}{n}z^{n},$
is algebraic over $\mathbb{Q}(x)$ if and only if $\alpha\in\mathbb{Q}$. We will assume by contradiction that a root $\alpha$ is irrational, although its annihilating polynomial for almost all prime numbers $p$ splits into linear factors.
The next ingredient is Hermite-Padй approximation. The ideas of this method were first introduced by Hermite in [Her73] to prove the transcendence of $e$, and then extended by him and Padй, his student [Her93, Pad92, Pad94]. It produces for a finite number of power series, a list of polynomials of prescribed degrees, such that the sum of the power series, weighted by the polynomials, has highest possible order. More precisely, let $K$ be a field. Let $f_{1}(z),\ldots,f_{m}(z)\in K[\![z]\!]$ be a tuple of power series. Let $n_{1},\ldots,n_{m}$ be non-negative integers and let $P_{1}(z),\ldots,P_{m}(z)\in K[z]$ be polynomials of degree $n_{1},\ldots,n_{m}$ respectively. If the power series
$g(z)\coloneqq P_{1}(z)f_{1}(z)+\cdots+P_{m}(z)f_{m}(z)\in K[\![z]\!]$
has order greater than or equal to $n_{1}+\cdots+n_{m}+m-1$ at 0, then the tuple $(P_{1}(z),\ldots,P_{m}(z))$ is called a Hermite-Padй approximant to $(f_{1}(z),\ldots,f_{m}(z))$ of type $(n_{1},\ldots,n_{m})$, and $g(z)$ is called the remainder of the approximation. For further exposition of the concept, see [Mah68, Jag64, Jag64a].
In the following, a well-known formula for the Hermite-Padй approximants of the consecutive powers $(1-z)^{i\alpha}$ with $\alpha\not\in\mathbb{Q}$ for $1\leq i\leq 2M+1$, with coefficients of uniform degree $N$ for any integer values of $M$ and $N$ is exploited. It was introduced by Padй in [Pad01] and studied and extended by Jager, Mahler, and G. Chudnovsky [Jag64a, Mah68, Chu83], but could already be deduced from Hermite’s work [Her73]. The result also includes an explicit expression for the first non-vanishing term in the approximation.
It is clear that this first non-vanishing term of the approximation will be an algebraic number. In the following one considers a certain multiple of this coefficient. One can estimate its norm, and its denominator by two different estimates, coming from either side of the equation for the Hermite-Padй approximation, depending on $M$ and $N$. The bound on the denominator uses the fact that if $\alpha$ reduces to an integer modulo a prime $p$, certain binomial coefficients involving $\alpha$ have non-negative $p$-adic valuation. From there, Chudnovsky and Chudnovsky use an asymptotic argument, that for sufficiently large $M$ and $N$, the bounds contradict the trivial inequality that the norm of a non-zero algebraic integer is greater or equal to one.
We will argue along the same lines, but we will not work asymptotically, but find upper bounds on the values of $M$ and $N$ for the contradiction to occur. In addition, we use that the argument of the Chudnovsky brothers relies on ruling out prime divisors of the denominators of a finite set of binomial coefficients, in whose denominators a priori only prime factors to a bound $\sigma$ depending on $M$ and $N$ may appear. This allows us to conclude if we assume that $R(w)$ factors into linear factors modulo $p$ for all primes $p$ up to $\sigma$.
For the rest of the section we fix an irrational number $\alpha$ with annihilating polynomial $R(w)=r_{n}w^{n}+\cdots+r_{1}w+r_{0}$. We set $\Delta\coloneqq|r_{n}|$, and pick $B\in\mathbb{R}$ to be a bound on the maximal modulus of a root of $R(w)$. We let $L$ be the splitting field of $R$. The denominator $\operatorname{den}(\gamma)$ of $\gamma\in L$ is the smallest positive integer $a\in\mathbb{N}_{>0}$, such that $a\gamma\in\mathcal{O}_{L}$, the ring of integers of $L$. In particular, we have $\operatorname{den}(\alpha)|\Delta$. Let $M,N$ be two positive integers and define $\sigma=\sigma(M,N)\coloneqq(2M+1)N+2M$. We will assume that $R(w)$ splits into linear factors for all primes up to $\sigma$. The proof of Theorem 1.2 is equivalent to showing that when choosing $M\coloneqq\left\lceil 2.826\cdot r_{n}^{3}\cdot\delta(\Delta)\right\rceil$ and $N\coloneqq 6.076\cdot BM$ the above assumptions lead to a contradiction. We will accomplish this in Proposition 4.12, contradicting Corollary 4.10.
4.1 Hermite-Padй Approximants to Binomial Powers and Estimates on Their Remainder
The explicit formula for Hermite-Padй approximants to the consecutive powers $(1-z)^{i\alpha}$ for $1\leq i\leq 2M+1$, with coefficients of uniform degree $N$ over $L$ reads
$\sum\limits_{i=1}^{2M+1}P_{i}(z)(1-z)^{(i-1)\alpha}=g(z)=\frac{N!^{2M+1}}{\sigma!}z^{\sigma}+O(z^{\sigma+1}),$ (4)
where $O(z^{\sigma+1})$ is a power series in $z^{\sigma+1}\mathbb{Q}[\![z]\!]$ and where
$P_{i}(z)\coloneqq\sum\limits_{h=0}^{N}p_{i,h}z^{h}\quad\text{with}\quad p_{i,h}\coloneqq\binom{N}{h}\left(\prod\limits_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{2M+1}\binom{(j-i)\alpha+N-h-1}{N}\right)^{-1}$ (5)
for all $1\leq i\leq 2M+1$ and $0\leq h\leq N$.
Note that the assumption on $\alpha$ not being a rational number is necessary for the expressions $p_{i,h}$ for large $N$ to be defined. At the same time for rational $\alpha$ the function $x^{\alpha}$ is algebraic, and there exist $M$ and a type, such that the remainder of the Hermite-Padй approximation vanishes.
In particular we know that the first nonzero coefficient in equation (4) is the coefficient of $z^{\sigma},$ which we will denote by $g_{\sigma}$. We obtain a second expression for it, by expanding the left hand side of (4):
$g_{\sigma}\coloneqq\sum\limits_{\mathclap{\begin{subarray}{c}1\leq i\leq 2M+1\\ 0\leq h\leq N\end{subarray}}}\ (-1)^{\sigma-h}\binom{(i-1)\alpha}{\sigma-h}p_{i,h}=\frac{N!^{2M+1}}{\sigma!}.$ (6)
In the following we will make use of the fact that the norm of a non-zero algebraic integer is at least $1$. More precisely, if $L$ is a number field of degree $d$ over $\mathbb{Q}$, then for any $\gamma\in L$ we have
$|\operatorname{den}(\gamma)^{d}\operatorname{Norm}_{L/\mathbb{Q}}(\gamma)|\geq 1.$ (7)
We will apply this inequality to $g_{\sigma}\cdot\Omega$, where $\Omega=\Omega_{M,N}$ is given by
$\Omega\coloneqq\left(\prod\limits_{k=1}^{M}\omega_{k}\right)\cdot\left(\prod\limits_{k=1}^{2M}\omega_{k}\right)\in L,\quad\text{with}\quad\omega_{k}\coloneqq\binom{k\alpha+N}{N}(k\alpha)\binom{-k\alpha+N}{N}$, (8)
to obtain the announced contradiction.
Remark 4.1.
Our definition of $\Omega$ differs slightly from the quantity called $\Omega$ in the Chudnovsky brothers’ article [CC85] because we believe that their line of argument, that we essentially followed and reproduced, does not hold with their choice. More precisely, the factors $(\pm k\alpha+N)$ do not appear in their definition of $\Omega$, hence for any $i\in\{1,\dots,2M+1\}$, the product $\Omega p_{i,N}$ simplifies to a polynomial in $\alpha$ with rational coefficients, times some factors of the form $(k\alpha+N)^{-1}$, for $k\in\mathbb{N}_{>0}$. The reasoning on the primes appearing in the denominator of this product does not hold anymore in this case (see the proof of Proposition 4.2, Equations (10) and (11)).
4.1.1 Bounding the Denominator
In this section we will prove the following bound on the denominator of $\Omega g_{\sigma}$.
Proposition 4.2.
The denominator of $\Omega g_{\sigma}$ satisfies
$\operatorname{den}(\Omega g_{\sigma})\leq C_{0}^{(2M-1)N}\Delta^{6MN+5M+N}\delta^{6MN+N+2M},$
where $C_{0}$ satisfies $\operatorname{lcm}(1,\dotsc,N)\leq C_{0}^{N}$ and $\delta=\prod\limits_{p|\Delta}p^{1/(p-1)}$.
Remark 4.3.
The existence of such a real number $C_{0}$ can be deduced from bounds on the Chebyshev function $\psi(x)=\log(\operatorname{lcm}(1,\dots,\lfloor x\rfloor))$. A possible $C_{0}$ is $\sqrt[113]{C}<2.826$, where $C$ is an explicit 51-digit number, see [RS62, Thm. 12] and OEIS A206431.
The following lemma is the key ingredient in the proof of Proposition 4.2, and as such, also the key ingredient to reducing Kronecker’s Theorem to a statement about a finite number of primes. We denote by $v_{p}(\cdot)$ the $p$-adic valuation.
Lemma 4.4.
Let $k,s\in\mathbb{Z}$, $r\in\mathbb{N}$, and let $p$ be a prime number. Let $\alpha$ be given as the zero of a polynomial $R(w)\in\mathbb{Z}[w]$ with leading coefficient $\Delta$.
  1.  (1)
    If $p\not\hskip 2.5pt\mid\Delta$ and $R(w)\bmod p$ splits completely in $\mathbb{F}_{p}[w]$, then $p\not\hskip 2.5pt\mid\operatorname{den}\left(\binom{k\alpha+s}{r}\right)$.
  2.  (2)
    If $p\not\hskip 2.5pt\mid\Delta$ and $p|\operatorname{den}\left(\binom{k\alpha+s}{r}\right)$, then $p<r$ and $v_{p}\left(\binom{k\alpha+s}{r}\right)\geq-\left\lfloor\frac{r}{p-1}\right\rfloor$.
  3.  (3)
    If $v_{p}(\Delta)=v$, then $v_{p}\left(\binom{k\alpha+s}{r}\right)\geq-\left(vr+\left\lfloor\frac{r}{p-1}\right\rfloor\right)$.
  4.  (4)
    If $R(w)\bmod p$ splits completely in $\mathbb{F}_{p}[w]$ for all primes $p\leq r$ not dividing $\Delta$, then $\operatorname{den}\left(\binom{k\alpha+s}{r}\right)$ divides $\Delta^{r}\prod_{p\mid\Delta}p^{\left\lfloor\frac{r}{p-1}\right\rfloor}.$
Доказательство.
Let $L/\mathbb{Q}$ be the splitting field of $R(w)$, and $\mathfrak{p}$ a prime ideal above $p$.
If $R(w)\bmod p$ splits in $\mathbb{F}_{p}[w]$, then there is a rational integer $a$ such that $\alpha\equiv a\bmod\mathfrak{p}$ hence $\binom{k\alpha+s}{r}\equiv\binom{ka+s}{r}\bmod\mathfrak{p}$ and the latter is a rational integer. Hence $v_{\mathfrak{p}}\left(\binom{k\alpha+s}{r}\right)\geq 0$. Denominators are rational integers and $\ell\in\mathbb{Z}$ belongs to $\mathfrak{p}$ if and only if $p|\ell$. Thus $p$ does not divide the denominator of $\binom{k\alpha+s}{r}$.
The second statement follows from the fact that $\binom{k\alpha+s}{r}=(k\alpha+s)_{r}/r!$ where $(u)_{r}\coloneqq u(u-1)\cdots(u-r+1)$ denotes the falling factorial. Thus the denominator of $\binom{k\alpha+s}{r}$ is composed of a product of primes dividing either $r!$ or the denominator of $\alpha$. Since $\operatorname{den}(\alpha)|\Delta$, we notice that $p\not\hskip 2.5pt\mid\operatorname{den}(\alpha)$ here. Primes dividing $r!$ are all primes smaller than $r$. Moreover it is a standard fact that $v_{p}(r!)\leq\frac{r}{p-1}$, hence the result.
For the third statement note that $p^{v}(k\alpha+s-\ell)$ has positive $p$-adic valuation for all integers $\ell$. Thus $v_{p}\left(p^{vr}(k\alpha+s)_{r}\right)\geq 0$ and we conclude as for the second statement.
The fourth assertion follows in a straightforward way from the previous three ones. ∎
Remark 4.5.
  1. (1)
    The proof of Lemma 4.4, is independent of the choice of the prime ideal $\mathfrak{p}$ over $p$ since $L$ is Galois, as the splitting field of a polynomial.
  2. (2)
    In the setting of the second assertion, we know from the contrapositive of the first assertion that $R(w)\bmod p$ does not split completely in $\mathbb{F}_{p}[w]$.
Let us recall some binomial identities which we will use in the following and can be proven by just rearranging factors. For any $x\in\mathbb{C}$, and $\ell,m\in\mathbb{N}$, with $\ell\geq m$ we have
$\binom{x}{\ell}\binom{\ell}{m}=\binom{x}{m}\binom{x-m}{\ell-m}=\binom{x}{\ell-m}\binom{x-\ell+m}{m}$. (9)
Proof of Proposition 4.2.
Let $M,N\in\mathbb{N}_{>0}$. For $1\leq i\leq 2M+1$, and $0\leq h\leq N$ consider the coefficients $p_{i,h}$ given in (5). In accordance with the middle expression of (6), we are aiming to bound $\operatorname{lcm}_{i,h}\left(\operatorname{den}\left(\Omega p_{i,h}\binom{(i-1)\alpha}{\sigma-h}\right)\right)$.
Let us start by rewriting the factors appearing in $p_{i,h}$. By applying the first equality of (9), we get
$\binom{N}{h}\binom{(j-i)\alpha+N-h-1}{N}=(-1)^{h}\binom{(j-i)\alpha+N-h-1}{N-h}\binom{(i-j)\alpha+h}{h}\eqqcolon q_{j-i,h}.$
Consequently we may write
$p_{i,h}=\binom{N}{h}^{2M+1}\cdot\prod_{k=1}^{i-1}q_{-k,h}^{-1}\prod_{k=1}^{2M+1-i}q_{k,h}^{-1}.$
Let us first treat the case $1\leq i\leq M$. We match the factors $\omega_{k}$ of $\Omega$ with the factors of $p_{i,h}$. Moreover, we note that $\binom{-k\alpha+N-h}{N-h}(k\alpha)=-(-k\alpha+N-h)\binom{-k\alpha+N-h-1}{N-h}$ and, using the right hand equality of (9), we obtain
$(-1)^{h}q_{-k,h}^{-1}\omega_{k}=\frac{\binom{k\alpha+N}{N}}{\binom{k\alpha+h}{h}}\cdot\frac{\binom{-k\alpha+N}{N}(k\alpha)}{\binom{-k\alpha+N-h-1}{N-h}}=\frac{\binom{k\alpha+N}{N-h}}{\binom{N}{h}}\cdot\frac{-\binom{-k\alpha+N}{h}(-k\alpha+N-h)}{\binom{N}{h}}$, (10)
where $k<i\leq M$. Analogously
$(-1)^{h}q_{k,h}^{-1}\omega_{k}=\frac{\binom{-k\alpha+N}{N-h}}{\binom{N}{h}}\cdot\frac{\binom{k\alpha+N}{h}(k\alpha+N-h)}{\binom{N}{h}}$, (11)
where $k<2M+1-i\leq 2M$.
In the case $M+1\leq i\leq 2M+1$ one proceeds analogously, where the matching of factors from $\Omega$ and $p_{i,h}$ is adapted accordingly.
In total, $\Omega p_{i,h}$ is of the form $\binom{N}{h}^{-2M+1}b,$ where $b$ is a product of binomial coefficients of the form $\binom{k\alpha+j}{h}$ and linear factors of the form $(k\alpha+j)$, consisting of the factors appearing in the numerators of the right-hand side of (10) and (11), and the $M$ ‘‘unused’’ factors $\omega_{k}$ for $i\leq k\leq M$ and $2M+2-i\leq k\leq 2M$. The degree of $b$ in $\alpha$ is given by $2M(N+1)+(2N+1)M=4MN+3M$. Indeed, in the $2M$ factors coming from (10) and (11) the degree is $N+1$ each and the remaining $M$ factors $\omega_{k}$ are of degree $2N+1$ each. Moreover, using that $\lfloor(N-h)/s\rfloor+\lfloor h/s\rfloor\leq\lfloor N/s\rfloor$ for any $s\in\mathbb{N}_{>0}$, we obtain from Lemma 4.4 (4) that $\operatorname{den}(b)$ divides
$\Delta^{4MN+3M}\prod_{p|\Delta}\left(p^{\left\lfloor\frac{N}{p-1}\right\rfloor}\right)^{4M}\left(\operatorname{lcm}_{h\in\{0,\ldots,N\}}\binom{N}{h}\right)^{2M-1}.$
Moreover, $\operatorname{den}\left(\binom{(i-1)\alpha}{\sigma-h}\right)$ divides the expression $\Delta^{\sigma-h}\prod_{p|\Delta}\left(p^{\left\lfloor\frac{\sigma-h}{p-1}\right\rfloor}\right),$ by Lemma 4.4 (4), which in turn divides $\Delta^{\sigma}\prod_{p|\Delta}\left(p^{\left\lfloor\frac{\sigma}{p-1}\right\rfloor}\right).$
Putting things together, for any pair $(i,h)$ we have that $\operatorname{den}(\Omega p_{i,h}\binom{(i-1)\alpha}{\sigma-h})$ divides
$\Delta^{6MN+5M+N}\prod_{p|\Delta}\left(p^{\left\lfloor\frac{6MN+N+2M}{p-1}\right\rfloor}\right)\left(\operatorname{lcm}_{h\in\{0,\ldots,N\}}\binom{N}{h}\right)^{2M-1}.$
Using that $\operatorname{lcm}_{h\in\{0,\ldots,N\}}\binom{N}{h}\leq\operatorname{lcm}\{1,\ldots,N\}\leq C_{0}^{N}$ we get
$\operatorname{den}(\Omega g)\leq\Delta^{6MN+5M+N}\delta^{6MN+N+2M}C_{0}^{(2M-1)N},$
as announced. ∎
4.1.2 Bounding the Norm
In this section we want to obtain a bound on $|\operatorname{Norm}(\Omega g_{\sigma})|,$ where $g_{\sigma}=\frac{N!^{2M+1}}{\sigma!}.$ Our computations will involve the sum of all integers from $1$ to $2M$, with those up to $M$ counted twice, as well as the sum of squares of these numbers. Thus we define the corresponding polynomials $u_{1}(M)$ and $u_{2}(M)$ to simplify our expressions
$u_{1}(M)\coloneqq\frac{5}{2}M^{2}+\frac{3}{2}M\hskip 14.22636pt\text{ and }\hskip 14.22636ptu_{2}(M)\coloneqq 3M^{3}+\cfrac{5}{2}M^{2}+\cfrac{1}{2}M.$
Also, we will make use of the following convenient fact about computing norms in a Galois extension.
Remark 4.6.
Let us remind that if $L/\mathbb{Q}(\alpha)/\mathbb{Q}$ is a Galois extension, and $\phi(\alpha)\in\mathbb{Q}(\alpha)$ is a rational expression in $\alpha$, then $\operatorname{Norm}_{L/\mathbb{Q}}(\phi(\alpha))\in\mathbb{Q}$ can be computed as follows, up to its sign. Let $\pi_{\alpha}(x)\in\mathbb{Q}[x]$, be the monic minimal polynomial of $\alpha$ of degree $e$ dividing $d\coloneqq[L:\mathbb{Q}]$, then $\operatorname{Norm}_{L/\mathbb{Q}}(\phi(\alpha))=\prod_{\tilde{\alpha}}\phi(\tilde{\alpha})^{d/e}$, where the product ranges over all roots $\tilde{\alpha}$ of $\pi_{\alpha}(x)$. In particular if $r\in\mathbb{Q}$, then $\operatorname{Norm}_{L/\mathbb{Q}}(r)=r^{d}$.
Proposition 4.7.
Let $B$ be an upper bound on the modulus of roots of $R(w)$, let $M\in\mathbb{N}$ and $N\geq 4MB$. Then
$|\operatorname{Norm}(\Omega g_{\sigma})|^{1/d}<\frac{(2e^{2})^{M}\cdot\exp\left(\pi u_{1}(M)B+4\log\left(\frac{4}{3}\right)\frac{u_{2}(M)}{N}B^{2}+\frac{M+1}{6N}\right)}{(2M+1)^{(2M+1)N}\cdot N^{M}\cdot(2M+1)^{2M+1/2}\cdot\pi^{2M}},$
where $B\in\mathbb{N}$ is a bound on the maximal modulus of all roots of a given annihilating polynomial $R(w)$ of $\alpha$.
To obtain bounds on the factorials, we borrow the following variant of Stirling’s formula from [Rob55]: For $k\in\mathbb{N}_{>0}$ we have
$\left(\frac{k}{e}\right)^{k}\cdot\sqrt{2\pi k}\cdot\exp\left(\frac{1}{12k+1}\right)<k!<\left(\frac{k}{e}\right)^{k}\cdot\sqrt{2\pi k}\cdot\exp\left(\frac{1}{12k}\right).$ (12)
We deduce the following bound on $g_{\sigma}$ by routine estimates, proving half of Proposition 4.7.
Lemma 4.8.
Let $M,N\in\mathbb{N}$ and write $\sigma=(2M+1)N+2M$. Then
$\frac{N!^{2M+1}}{\sigma!}<(2M+1)^{-(2M+1)N}N^{-M}(2M+1)^{-2M-1/2}\left(2\pi e^{2}\right)^{M}\exp\left(\frac{M+1}{6N}\right).$
The second part of the proof of Proposition 4.7 is using Euler’s product formula for the sine function, stating that for $z\in\mathbb{C}$ we have
$\sin(z)=z\prod\limits_{j\geq 0}\left(1-\frac{z^{2}}{j^{2}\pi^{2}}\right).$ (13)
Lemma 4.9.
Assuming that $N\geq 4MB$ we have
$|\operatorname{Norm}(\Omega)|^{1/d}\leq\frac{1}{\pi^{3M}}\cdot\exp\left(\pi Bu_{1}(M)+4\log(4/3)\frac{B^{2}}{N}u_{2}(M)\right).$
Доказательство.
For a fixed integer $k\in\{1,\ldots,2M\}$ we can rewrite the corresponding factor appearing in $\Omega$ using (13) as
$\binom{k\alpha+N}{N}(k\alpha)\binom{-k\alpha+N}{N}=(k\alpha)\prod\limits_{j=1}^{N}\left(1-\frac{k^{2}\alpha^{2}}{j^{2}}\right)=\frac{\sin(k\pi\alpha)}{\pi}\ \prod_{\mathclap{j\geq N+1}}\ \left(1-\frac{k^{2}\alpha^{2}}{j^{2}}\right)^{-1}.$
We now wish to bound the sine by the exponential function, and estimate the remaining product. For any $z\in\mathbb{C}$, we have $|\sin(z)|\leq\exp(|z|)$ hence $|\sin(k\pi\alpha)|\leq\exp(k\pi B)$, as $B$ is a bound on the modulus of $\alpha$.
For the remaining product, let us now write $x_{j}\coloneqq\frac{k^{2}\alpha^{2}}{j^{2}}$. By our assumption on $N$, we have that $|x_{j}|\leq\frac{1}{4}$. We then make use of the fact that the function $f(x)\coloneqq(1-x)^{1/x}$ is strictly decreasing on the interval $(0,1)$ and deduce that $0<|x_{j}|\leq 1/4$ implies $f(1/4)\leq f(|x_{j}|)$, which is, after applying the logarithm on both sides, equivalent to $\log(1-|x_{j}|)\leq 4\log(4/3)\cdot|x_{j}|$. We obtain the bound
$\left|\prod_{j\geq N+1}(1-x_{j})^{-1}\right|$ $\leq\prod_{j\geq N+1}(1-|x_{j}|)^{-1}=\exp\left(\sum_{j\geq N+1}-\log(1-|x_{j}|)\right)$
$\leq\exp\left(4\log\left(\frac{4}{3}\right)\cdot\sum_{\mathclap{j\geq N+1}}|x_{j}|\right)\leq\exp\left(4\log\left(\frac{4}{3}\right)\cdot\frac{k^{2}B^{2}}{N}\right),$
where the last inequality is justified by $|x_{j}|\leq(kB)^{2}/j^{2}$, and $\sum_{j\geq N+1}\frac{1}{j^{2}}\leq\frac{1}{N}$.
We take the product of the bounds obtained for each individual $k$ for $k\in\{1,\dots,2M\}$ and, again, for $k\in\{1,\dots,M\}$ to obtain
$|\Omega|^{1/d}\leq\frac{1}{\pi^{3M}}\cdot\exp\left(\pi Bu_{1}(M)+4\log(4/3)\frac{B^{2}}{N}u_{2}(M)\right).$
All the estimates above are equally valid when replacing $\alpha$ by any root of $R(w)$, so in particular for the conjugates of $\alpha$. Thus the estimates also hold for any conjugate of $\Omega$, and, by Remark 4.6, we conclude. ∎
Putting together Lemmata 4.8 and 4.9 finishes the proof of Proposition 4.7.
4.2 A Bound on the Number of Primes
Finally, combining the results of Propositions 4.2 and 4.7 we get an upper bound on the integer $|\operatorname{den}(\Omega g_{\sigma})^{d}\operatorname{Norm}_{L/\mathbb{Q}}(\Omega g_{\sigma})|$. To prepare for further estimates, we split it into three parts and set
$\begin{array}[]{cl}X(M)&\coloneqq\cfrac{C_{0}^{2M-1}\Delta^{6M+1}\delta^{6M+1}}{(2M+1)^{2M+1}}=\cfrac{1}{C_{0}^{2}\Delta^{2}\delta^{2}}\cdot\left(\cfrac{C_{0}\Delta^{3}\delta^{3}}{2M+1}\right)^{2M+1}\\ Y(M,N)&\coloneqq\exp\left(\pi Bu_{1}(M)+4\log\left(\cfrac{4}{3}\right)B^{2}\cfrac{u_{2}(M)}{N}\right)\cdot(2M+1)^{-2M-1/2}\cdot\left(\cfrac{2e^{2}\Delta^{5}\delta^{2}}{\pi^{2}}\right)^{M}\\ Z(M,N)&\coloneqq\left(\cfrac{1}{N}\right)^{M}\cdot\exp\left(\cfrac{M+1}{6N}\right).\end{array}$
Of course, these quantities also depend on $\Delta$, which is suppressed in the notation. Moreover, once we fix $M$ to be a function of $\Delta$, the notation becomes particularly misleading. However, in the following calculations it will be convenient to view these quantities as functions of $M$ and $N$.
Corollary 4.10.
Let $M\in\mathbb{N}$ and $N\geq 4MB.$ Then
$0<|\operatorname{den}(\Omega g_{\sigma})\operatorname{Norm}_{L/\mathbb{Q}}(\Omega g_{\sigma})^{\frac{1}{d}}|<X(M)^{N}\cdot Y(M,N)\cdot Z(M,N).$ (14)
From the asymptotics of these bounds, one can easily deduce the existence of parameters $M$ and $N$, such that $|\operatorname{den}(\Omega g_{\sigma})\operatorname{Norm}_{L/\mathbb{Q}}(\Omega g_{\sigma})^{\frac{1}{d}}|<1$. For large enough $M$ compared to $\Delta$, the expression $X(M)$ is strictly smaller than $1$. Moreover, if we choose $N$ large enough with respect to $M$, we have $X(M)^{N}<Y(M,N)^{-1}$, as the right-hand side is independent of $N$. Finally, for $N>M>3$, clearly $Z(M,N)<1$. Thus, for this choice of $M,N$ we find the desired contradiction to the assumption that $\alpha$ was irrational, and so we have proved Kronecker’s Theorem. This argumentation was carried out by the Chudnovsky brothers in [CC85]. The rest of this subsection is devoted to showing that our explicit choices of $M$ and $N$ suffice for the contradiction to occur.
4.2.1 Choice of the Number of Functions to Approximate
Let us first justify our choice of $M$.
Lemma 4.11.
For any $M>\frac{1}{2}(C_{0}\Delta^{3}\delta^{3}-1)$ we have $X(M)<1$. Moreover, for $M>C_{0}\Delta^{3}\delta^{3}$ we have $X(M)<2^{-(2M+1)}(C_{0}\Delta\delta)^{-2}$.
Доказательство.
The proof consists of a straightforward computation. ∎
The choices of $M$ and $N$ are dependent on each other in the following way: the function $X(M)$ is decreasing with $M$, while simultaneously $Y(M,N)$ increases. Thus, the first value of $N$, such that $X(M)^{N}Y(M,N)Z(M,N)<1$ varies in a non-obvious way with $M$. To obtain tight bounds one should minimize $\sigma=(2M+1)N+2M$ under the constraint $X(M)^{N}Y(M,N)Z(M,N)<1$. We carry out no such computations, but are contempt with $M=\lceil C_{0}\Delta^{3}\delta^{3}\rceil$. Naïve estimates and computations suggest that picking $M\approx\lceil 0.806\cdot C_{0}\Delta^{3}\delta^{3}\rceil$ yields slightly more optimal values for $\sigma$.
4.2.2 Choice of the Type in the Approximation
The goal of this section is to prove that for the explicit values of $M$ and $N$ given in Theorem 1.2 the right-hand side of (14) is bounded from above by $1$. We assume $M=\lceil C_{0}\Delta^{3}\delta^{3}\rceil$ to be fixed. In Proposition 4.7, we assume $N$ to depend at least linearly on $B$ and $M$. Let us set $N\coloneqq ABM$ for some constant $A\geq 4$ to be determined later. With these choices of $M$ and $N$ the right-hand side of (14) is a function in $\Delta$. However, to simplify computations we will treat it as a function in $M=\lceil C_{0}\Delta^{3}\delta^{3}\rceil$.
Proposition 4.12.
For $M=\lceil C_{0}\Delta^{3}\delta^{3}\rceil$ and $N=ABM$ with $A\geq 4$, we have
$X(M)^{N}Y(M,N)Z(M,N)<\exp\left(T(M)\right),$ (15)
where $T(M)\coloneqq c_{2}M^{2}+\tilde{c}_{1}M\log(M)+c_{1}M+\tilde{c}_{0}\log(M)+c_{0}$ with
$c_{2}\coloneqq\left(-2A\log(2)+\frac{5\pi}{2}+\frac{12\log(4/3)}{A}\right)B,\qquad\tilde{c}_{1}\coloneqq-1-\frac{2}{3}AB$
$c_{1}\coloneqq\left(-A\left(\frac{2}{3}\log\left(\frac{2C_{0}^{2}}{3}\right)+\log(2)\right)+\frac{3\pi}{2}+\frac{10\log\left(\frac{4}{3}\right)}{A}\right)B,$
$\tilde{c}_{0}\coloneqq-\frac{1}{2},\qquad c_{0}\coloneqq\frac{1}{18}+\frac{2\log\left(\frac{4}{3}\right)B}{A}.$
For $A>6.076$ and all $B\geq 1$ and $\Delta\geq 1$ the expression $T(M)$ is negative and thus
$X(M)^{N}Y(M,N)Z(M,N)<1.$
Доказательство.
We estimate crudely $\frac{2e^{2}\Delta^{5}\delta^{2}}{N\pi^{2}}<M$. Moreover, by Lemma 4.11, $X(M)<(C_{0}\Delta\delta)^{-2}\cdot 2^{-(2M+1)}$. We have $\log(C_{0}^{2}\Delta^{2}\delta^{2})=-\frac{4}{3}\log(C_{0})-\frac{2}{3}\log(C_{0}\Delta^{3}\delta^{3})$, and $\log(C_{0}\Delta^{3}\delta^{3})\leq-\log(M-1)\leq-\log(M)-\log(\frac{2}{3})$ by factoring $M$ in the $\log$ and using that $M\geq 3$, hence $\log(C_{0}^{2}\Delta^{2}\delta^{2})\leq-\frac{2}{3}\log(\frac{2C_{0}^{2}}{3})-\frac{2}{3}\log(M)$. We further use $\log(2M+1)>\log(M)$ and bound $\frac{M+1}{6N}<\frac{1}{18}$. The rest of the proof of the first part is obtained from the definitions of $X(M)$ and $Y(M,N)$ by straightforward computations.
Finding $A$ such that $c_{2}$ is negative amounts to solving a quadratic equation in $A$, yielding $c_{2}<0$ for $A>6.076$. Similarly, we find $c_{1}<0$ for $A\geq 3$. As $M\geq 3$, we note that $\frac{1}{2}\log(M)+\frac{1}{18}<0$, and $3\log(3)\leq M\log(M)$ hence $\frac{2}{3}AM\log(M)+\frac{2\log\left(\frac{4}{3}\right)}{A}<0$ for $A\geq 1$. So $\tilde{c}_{1}M\log(M)+\tilde{c}_{0}\log(M)+c_{0}<0$. Altogether, we have proved that $T(M)<c_{2}M^{2}$, assuming that $A\geq 4$, and this concludes the proof. ∎
5 Algorithm and Complexity
The bounds we present in Theorem 1.3 allow us, as was our purpose, to solve algorithmically the problem of deciding the algebraicity of solutions of equations (2). We present here an algorithm, Algorithm 3, solving this problem and we estimate its complexity. Each of its steps is studied in what follows, and the full algorithm’s complexity estimate is analyzed in Section 5.5. All complexities are stated in number of bit operations.
Throughout this section, we consider our input rational function $u(x)$ to be of normal form $\frac{a(x)}{b(x)}$ where $a(x)$ and $b(x)$ have integer coefficients, are primitive, coprime, and $\deg a(x)<\deg b(x)$. The primitivity condition can be made without loss of generality, as already discussed in the proof of Theorem 1.3 in Section 2. This comes from the fact that if $y(x)$ satisfies $y^{\prime}(x)=u(x)y(x)$, then for any $c\in\mathbb{Q}$, the function $\tilde{y}(x)\coloneqq y(x)^{c}$ satisfies $\tilde{y}^{\prime}(x)=c\cdot u(x)\tilde{y}(x)$, and $y(x)$ is algebraic if and only if $\tilde{y}(x)$ is algebraic.
Our measure of complexity of a rational number will be an estimate of ‘‘its size on a computer’’, that we call its height. More precisely, if $q=a/b\in\mathbb{Q}$ with $a,b\in\mathbb{Z}\setminus\{0\}$ coprime, the height of $q$ is $h(q)\coloneqq\log_{2}(|ab|)$, and we set the height of 0 to be $1$. The height $h(f(x))$ of $f(x)\in\mathbb{Q}[x]$ is the maximum of the heights of its coefficients, and its multiplicative height is $H(f(x))\coloneqq 2^{h(f(x))}$. Remark that if $f(x)$ has integer coefficients, then $H(f(x))$ is the maximum of the modulus of its coefficients. The height of a rational function $u(x)=a(x)/b(x)$ with $a(x),b(x)\in\mathbb{Z}[x]$ primitive and coprime, is the maximum of $h(a(x))$ and $h(b(x))$.
The computation of the normal form, as described above, of $u(x)\in\mathbb{Q}(x)$ of height $h$ and degree n1n-1 in the numerator and $n$ in the denominator requires computing the greatest common divisor of $n+1$ numbers and the greatest common divisor of two polynomials. The computation of the greatest common divisor of two polynomials in $\mathbb{Q}[x]$ of degree at most $n$ and height $h$ can be made by performing $\tilde{O}(nh)$ bit operations [vzGG13, Corollary 11.9, p.325]. With an adaptation of the same algorithm we can compute the gcd of two integers of height at most $h$ in $\tilde{O}(h)$ bit operations. Hence applying it $n$ times allows to compute the greatest common divisor of $n+1$ integers of height at most $h$ in $\tilde{O}(nh)$ bit operations, thus computing the normal form of $u(x)$ can be made by performing $\tilde{O}(nh)$ bit operations.
Algorithm 3 Deciding algebraicity with $p$-curvatures.
[…]
1:if $\deg a(x)\geq\deg b(x)$ or $b(x)$ is not square-free then return Transcendental;
2:Compute the normal form of $a(x)/b(x)$;
[…]
6:while $p\leq\Delta$ do
7:  if $\Delta\bmod p=0$ then $\delta\leftarrow\delta\cdot p^{1/(p-1)}$
8:  else if $\texttt{pCurvature}(a(x),b(x),p)\neq 0$ then return Transcendental;
9:  $p\leftarrow\texttt{nextprime}(p)$
10:Compute $B$, $M\leftarrow\lceil 2.826\Delta^{3}t^{3}\rceil$, $N\leftarrow\lceil 6.076BM\rceil$;
[…]
12:while $p\leq\sigma$ do
13:  if $\texttt{pCurvature}(a(x),b(x),p)\neq 0$ then return Transcendental;
14:  $p\leftarrow\texttt{nextprime}(p)$;
15:return Algebraic;
Remark 5.1.
As explained in Remark 3.1, starting from $u(x)\in\overline{\mathbb{Q}}(x)$, we can check if $\tilde{R}(w)\coloneqq R(w)/\Delta$ is in $\mathbb{Q}[w]$ after step 4. If this is not the case we can immediately conclude that the output has to be Transcendental with Proposition 2.6. Otherwise, we take $\Delta$ to be the discriminant of the polynomial in $\mathbb{Z}[w]$ associated to $\tilde{R}(w)$.
5.1 Complexity of Rothstein-Trager Resultants
Let $u(x)=a(x)/b(x)\in\mathbb{Q}(x)$ be in normal form with $\deg a(x)<n\coloneqq\deg b(x)$, with the denominator $b(x)$ being squarefree, height bounded by $h$ and multiplicative height at most $H=2^{h}$.
Proposition 5.2.
The Rothstein-Trager resultant $R(w)\coloneqq\textnormal{res}_{x}(b(x),a(x)-w\cdot b^{\prime}(x))$ can be computed using $\tilde{O}(n^{2}h)$ bit operations.
Доказательство.
The coefficient of $x^{n-1}$ in $a(x)-w\cdot b^{\prime}(x)$ cannot vanish hence $\deg_{x}(a(x)-w\cdot b^{\prime}(x))=n-1$. Corollary 11.21 from [vzGG13] concludes. ∎
If we write $b(x)=\sum_{i=0}^{n}b_{i}x^{i}$ and $a(x)=\sum_{i=0}^{n-1}a_{i}x^{i}$ then the Rothstein-Trager resultant is the determinant of the $(2n-1)\times(2n-1)$ matrix
$\begin{pmatrix}b_{n}&&&&a_{n-1}-nwb_{n}&&&&\\ b_{n-1}&b_{n}&&&\vdots&a_{n-1}-nwb_{n}&&&\\ \vdots&b_{n-1}&\ddots&&\vdots&\vdots&\ddots&&\\ \vdots&\vdots&\ddots&b_{n}&a_{1}-2wb_{2}&\vdots&&\ddots&\\ b_{1}&\vdots&&b_{n-1}&a_{0}-wb_{1}&a_{1}-2wb_{2}&&&a_{n-1}-nwb_{n}\\ b_{0}&b_{1}&&\vdots&&a_{0}-wb_{1}&\ddots&&\vdots\\ &b_{0}&\ddots&\vdots&&&\ddots&\ddots&\vdots\\ &&\ddots&b_{1}&&&&\ddots&a_{1}-2wb_{2}\\ &&&b_{0}&&&&&a_{0}-wb_{1}\end{pmatrix}$. (16)
From the form of this Sylvester matrix, we see that the degree of $R(w)$ is at most $n$, and its coefficient of degree $n$ is $\operatorname{res}(b(x),b^{\prime}(x))$, up to a sign. The assumption that $b(x)$ is squarefree ensures $\deg(R(w))=n$. Concerning the height of $R(w)$ we prove the following estimate.
Proposition 5.3.
For any $0\leq k\leq n$ the coefficient $r_{k}$ of $w^{k}$ in $R(w)=\operatorname{res}_{x}(b(x),a(x)-w\cdot b^{\prime}(x))$ satisfies the inequality
$|r_{k}|\leq\binom{n}{k}6^{-k/2}H^{2n-1}(n+1)^{(n+k-1)/2}n^{n/2}(2n+1)^{k/2}.$
Remark 5.4.
The determinant is linear with respect to its columns, hence we can write the last columns of the matrix (16) each as the sum of a part with the coefficients of $a(x)$, and a part with the coefficients of $b^{\prime}(x)$, and factor $w$ in each column it appears. See Example 5.5 below for an illustration. The problem of bounding the height becomes bounding a determinant of a particular shape with integer coefficients and estimating the number of terms for each power of $w$ that appears.
Example 5.5.
In the case $n=2$, $b(x)=b_{0}+b_{1}x+b_{2}x^{2}$ and $a(x)=a_{0}+a_{1}x$, the polynomial $R(w)=\operatorname{res}_{x}(b(x),a(x)-w\cdot b^{\prime}(x))$ can be expanded as follows
$R(w)$ $=\begin{vmatrix}b_{2}&a_{1}-2wb_{2}&0\\ b_{1}&a_{0}-wb_{1}&a_{1}-2wb_{2}\\ b_{0}&0&a_{0}-wb_{1}\end{vmatrix}$
$=\begin{vmatrix}b_{2}&a_{1}&0\\ b_{1}&a_{0}&a_{1}\\ b_{0}&0&a_{0}\end{vmatrix}-w\left(\begin{vmatrix}b_{2}&2b_{2}&0\\ b_{1}&b_{1}&a_{1}\\ b_{0}&0&a_{0}\end{vmatrix}+\begin{vmatrix}b_{2}&a_{1}&0\\ b_{1}&a_{0}&2b_{2}\\ b_{0}&0&b_{1}\end{vmatrix}\right)+w^{2}\begin{vmatrix}b_{2}&2b_{2}&0\\ b_{1}&b_{1}&2b_{2}\\ b_{0}&0&b_{1}\end{vmatrix}$
The main ingredient for our estimates is the following lemma known as Hadamard’s inequality [vzGG13, Thm. 16.6, p. 477].
Lemma 5.6.
Let $M$ be a $n\times n$ matrix over $\mathbb{R}$, with columns $C_{1},\dots,C_{n}$, and coefficients bounded by $B>0$. Then
[…]
where $C||$ denotes the $2$-norm of the vector $C\in\mathbb{R}^{n}$.
Proof of Proposition 5.3..
Following the idea of using linearity presented in the Example 5.5, we can write each coefficient $r_{k}$ as a sum of determinants. Those determinants have a left block of n1n-1 columns with the coefficients of $b(x)$, and a right block of $n$ columns, each being either coefficients of $a(x)$ or coefficients of $b^{\prime}(x)$. Note that the power of $w$ in front of such a determinant is equal to the number of columns corresponding to coefficients of $b^{\prime}(x)$. We will use Hadamard’s inequality, Lemma 5.6, on each of these determinants, hence let us compute a sharper bound on the $2$-norms of each possible column. A column $C_{1}$ of coefficients of $b(x)$ has at most $n+1$ nonzero coefficients, each bounded by $H=\exp(h)$, hence $C_{1}||=\left(\sum\limits_{i=0}^{n}b_{i}^{2}\right)^{1/2}\leq(n+1)^{1/2}H$. Similarly, a column $C_{2}$ of coefficients of $a(x)$ has at most $n$ nonzero coefficients, each bounded by $H$, hence $C_{2}||\leq n^{1/2}H$. Finally, if $C_{3}$ is a column of coefficients of $b^{\prime}(x)$, then $C_{3}||\leq\left(\sum\limits_{i=1}^{n}(ib_{i})^{2}\right)^{1/2}\leq\left(\frac{n(n+1)(2n+1)}{6}\right)^{1/2}H$.
[…]
Remark 5.7.
Starting from $u(x)\coloneqq\frac{a(x)}{b(x)}\in\overline{\mathbb{Q}}(x)\setminus\mathbb{Q}(x)$, the computations are more costly. Indeed, in a number field $\mathbb{Q}(\alpha)=\mathbb{Q}[t]/(\pi_{\alpha}(t))$ with $\pi_{\alpha}(t)\in\mathbb{Q}[t]$ the minimal polynomial of $\alpha\in\overline{\mathbb{Q}}$, elements are polynomials of degree smaller than $\deg(\pi_{\alpha}(t))$, hence algebraic number multiplication has the complexity of modular polynomial multiplication.
5.2 Complexity Estimates for Finding Rational Roots of the Rothstein-Trager Resultant
In Section 3.1 we noted algebraicity of the solutions of (2) can be decided by Algorithm 1. We briefly investigate the complexity of this algorithm. For efficiently finding all the rational roots of a polynomial we have the following complexity [Bos+17, Prop. 21.22].
Theorem 5.8.
Let $R(w)\in\mathbb{Z}[w]$ be a squarefree, primitive polynomial of degree $n$ and height $h$. The computation of all rational roots of $R(w)$ can be performed, in $\tilde{O}(n^{2}h)$ bit operations.
Remark 5.9.
Von zur Gathen and Gerhard only give a probabilistic algorithm performing the task in the same expected complexity [vzGG13, Thm. 15.21]. The algorithm described in [Bos+17] uses a deterministic (potentially costly) algorithm to find roots of the polynomial modulo a deterministically defined small prime number, and then lifts them to characteristic zero. It improves a result by Loos [Loo83], who proposed a deterministic algorithm for finding rational roots in $\tilde{O}(n^{3}h)$ bit operations.
Using the corresponding algorithm the cost of computing the Rothstein-Trager resultant is comparable to finding rational roots, see Proposition 5.3. We obtain the following Corollary.
Corollary 5.10.
Given $u(x)\coloneqq\frac{a(x)}{b(x)}\in\mathbb{Q}(x)$ of degree $n$ and height $h$, deciding if all solutions of $y^{\prime}(x)=u(x)y(x)$ are algebraic can be done by performing $\tilde{O}(n^{2}h)$ bit operations.
5.3 Computation of the Bound on the Primes
Given $u(x)$, and having computed the Rothstein-Trager resultant $R(w)\in\mathbb{Z}[w]$, Theorem 1.2 gives an explicit bound $\sigma\in\mathbb{N}$ on the number of $p$-curvatures to check, depending on $\Delta$, the leading coefficient, and an upper bound $B$ on the modulus of all roots of $R(w)$.
An arbitrarily precise estimate of $B$ is possible without factoring completely the polynomial $R(w)$, for example by following ideas from an unpublished report of Schцnhage in which an algorithm to find the complex roots of a polynomial $R(w)\in\mathbb{C}[w]$ with arbitrary precision is presented [Sch82, Thm. 15.1].
Proposition 5.11 (Schцnhage).
Let $R(w)\in\mathbb{Z}[w]$ of degree $n$ and $\tau>0$. The computation of $B_{0}>0$ such that the maximum modulus $r$ of all roots of $R(w)$ satisfies $B_{0}e^{-\tau}<r<B_{0}e^{\tau}$ can be done in $O(n^{2}(\log(\frac{1}{\tau})+\log(\log(n)))\log(\frac{4}{\tau}))$ binary operations.
The bound $B$ we are looking for is $B_{0}e^{\tau}$. For estimating the complexity of Algorithm 3, we will see that simply taking $\tau=1/2$ suits us, hence the complexity of computing the bound $B$ is $O(n^{2})$ bit operations.
Before computing a $p$-curvature, we need to ensure that $p\not\hskip 2.5pt\mid\Delta$, hence we can find the prime factors of $\Delta$ step by step, without factoring $\Delta$. After checking all $p$-curvatures for $p\leq\Delta$, we can compute $\delta$, then $M$, NN and $\sigma$.
5.4 Checking the Vanishing of $p$-Curvatures
Our goal in this section is to explain how to efficiently check if the $p$-curvatures of equation (2) vanish. We recall that the $p$-curvature of equation $y^{\prime}(x)=u(x)y(x)$, is $u^{(p-1)}(x)+u(x)^{p}\bmod p$. Consider now $u(x)\in\mathbb{Q}(x)$ of degree $n\in\mathbb{N}$ and multiplicative height at most $H>0$.
To compute $p$-curvatures we rely on an algorithm due to Bostan and Schost [BS09] which is tailored specifically for first-order differential equations. By ‘‘computing’’ a $p$-curvature, we mean that we compute enough terms of its Taylor expansion in order to uniquely reconstruct it knowing a bound on the degrees of its numerator and denominator. Indeed, a rational function with numerator’s and denominator’s degrees at most $n$ is completely determined by $n$ and the first $2n$ coefficients of its Taylor series expansion.
Algorithm 4 pCurvature (Computing one $p$-curvature.)
Input: polynomials $a(x),b(x)\in\mathbb{Z}[x]$, a prime number $p$.
   Output: The $p$-curvature of $y^{\prime}(x)=\frac{a(x)}{b(x)}y(x)$.
1:if $p|\operatorname{res}_{x}(b(x),b^{\prime}(x))$ then return ‘Error, bad prime $p$’;
[…]
4:Compute $\bar{u}_{p-1}$;
5:for $i=2$ to $2n$ do compute $\bar{u}_{ip-1}$;
[…]
7:return $w(x)+v(x)$;
Proposition 5.12 (Bostan, Schost).
Let $u(x)\in\mathbb{Q}(x)$ of degree $n\in\mathbb{N}$, with coefficients bounded by $H>0$. For any prime number $p$ the computation of the $p$-curvature of $y^{\prime}(x)=u(x)y(x)$ can be performed using Algorithm 4 in $\tilde{O}(n\log(p)(n+\log(p)+\log(H)))$ bit operations. The computation of all $p$-curvatures of $y^{\prime}(x)=u(x)y(x)$ for $p\leq S$ can be done in $\tilde{O}(n^{2}(S+\log(H))+nS\log(S))$ bit operations, where we neglect factors $\log(n)$ and $\log(\log(p))$.
Доказательство.
Let us consider $a(x),b(x)\in\mathbb{Z}[x]$, coprime, and $y^{\prime}(x)=u(x)y(x)$ where $u(x)=\frac{a(x)}{b(x)}$. Let $p$ be a prime number not dividing $\operatorname{res}_{x}(b(x),b^{\prime}(x))$. We will denote the reduction of a rational number $c$, or a polynomial or rational function $u(x)$, modulo $p$, by $\bar{c}$, and $\bar{u}(x)$, respectively. Let us assume that $\bar{a}(x)$ and $\bar{b}(x)$ remain coprime in $\mathbb{F}_{p}[x]$, else we divide them by their greatest common divisor. The $p$-curvature $\psi(x)\coloneqq\bar{u}^{(p-1)}(x)+\bar{u}(x)^{p}$ is a rational function in $\mathbb{F}_{p}(x^{p})$. By linearity of taking the $1/p$-th power over $\mathbb{F}_{p}$, $\psi(x)^{1/p}=\psi(x^{1/p})$ is the sum of $\bar{u}(x)$ and $\bar{v}(x)\coloneqq(\bar{u}(x)^{(p-1)})^{1/p}$.
The computation of the expansion of $\bar{u}(x)$ by performing a Newton iteration can be done in $\tilde{O}(n\log(p))$ bit operations [vzGG13, Thm. 9.4, p. 260]. If we write $\bar{u}(x)=\bar{u}_{0}+\bar{u}_{1}x+\dots+\bar{u}_{i}\in\mathbb{F}_{p}$, then the Taylor expansion of $\bar{v}(x)$ is $\bar{u}_{p-1}+\bar{u}_{2p-1}x+\bar{u}_{3p-1}x^{2}+\dots$ The extraction of these specific coefficients can be done with an algorithm due to Fiduccia [Fid85], performing $\tilde{O}(n\log(p)(n+\log(p)))$ bit operations. This algorithm uses the following result proved by Fiduccia: if we write $f(x)\coloneqq x^{n-\operatorname{val}_{x}(b(x))}\bar{b}(\frac{1}{x})\in\mathbb{F}_{p}[x]$ for the characteristic polynomial of the sequence $(\bar{u}_{i})$, then for any $k\in\mathbb{N}_{\geq n}$, $\bar{u}_{k}=\varphi_{0,k}\bar{u}_{0}+\dots+\varphi_{n-1,k}\bar{u}_{n-1}$ where $\varphi_{i,k}\in\mathbb{F}_{p}$ are defined by $x^{k}=\varphi_{0,k}+\varphi_{1,k}x+\dots+\varphi_{n-1,k}x^{n-1}\bmod f(x)$. Knowing the first $2n$ terms $\bar{u}_{0},\dots,\bar{u}_{2n-1}$ of the Taylor expansion of $\bar{u}(x)$, the computation of the p1p-1-th term $\bar{u}_{p-1}$ amounts to compute a power of the image $\xi$ of $x$ in the ring $\mathbb{F}_{p}[x]/(f(x))$ and this can be made by binary powering in $\tilde{O}(n\log(p)^{2})$ bit operations. Now that we have computed $\xi^{p-1}$, it only takes one product in $\mathbb{F}_{p}[x]/(f(x))$ to compute $\xi^{p}$, and one more for each $\xi^{(i+1)p-1}$ knowing $\xi^{ip-1}$. In total computing the coefficients of $v(x)$ up to order $2n$ takes $\tilde{O}(n\log(p)(n+\log(p)))$ with this method.
The first two steps of Algorithm 4 consist almost only of Euclidean divisions, thus step 1 can be performed in $\tilde{O}((n^{2}+\log(p))\log(H))$, and step 2 in $\tilde{O}(n\log(p)\log(H))$. Based on the previous discussions, step 3 can be performed using $\tilde{O}(n\log(p))$ bit operations, steps 4 and 5 in respectively $\tilde{O}(n\log(p)^{2})$ and $\tilde{O}(n^{2}\log(p))$ bit operations. Then adding two polynomials of degree $2n$ takes $\tilde{O}(n\log(p))$ bit operations, thus the computation of one $p$-curvature can be made in $\tilde{O}(n\log(p)(n+\log(p)+\log(H)))$ bit operations as announced. The complexity of repeating this algorithm for all primes up to $S$ follows knowing the estimates $\sum_{p\leq S}\log(p)=O(S)$ and $\sum_{p\leq S}\log(p)^{2}=O(S\log(S))$. ∎
Remark 5.13.
An efficient algorithm to compute the characteristic polynomials of the $p$-curvatures for any order differential equations for all primes $p\leq S$ simultaneously in $\tilde{O}(S(\log(H)+n)n^{\omega+1})$ bit operations is given by Pagиs in [Pag21], where $\omega<2.3728596$ is an exponent for matrix multiplication. This is easily adapted to check the nilpotence of the $p$-curvatures in the same complexity. Checking nullity in such a complexity could prove crucial in the development of efficient algorithms for a potential effective version of the Grothendieck $p$-curvature conjecture for (classes of) higher order equations.
5.5 Complexity Estimate
Our algorithm to decide the nature of all solutions of $y^{\prime}(x)=u(x)y(x)$ by computing $p$-curvatures is presented above in Algorithm 3.
Let us justify that the $\delta$ from Theorem 1.3 is negligible in our estimates.
Lemma 5.14.
Let $\Delta\in\mathbb{N}$ and $\delta=\prod_{p|\Delta}p^{1/(p-1)}$, then $\delta=O(\log(\Delta)^{2})$.
Доказательство.
First we remark that if $\ell\in\mathbb{N}$ is such that $\ell\#\leq\Delta<(\ell+1)\#$, then $\delta(\Delta)\leq\delta(\ell\#)$ where $\ell\#\coloneqq\prod_{p\leq\ell}p$ is the primorial of $\ell$. This follows from the fact that $x^{1/(x-1)}$ is decreasing for $x\geq 2$. We will now estimate $\delta(\ell\#)$. In [RS62, Theorem 10], we find that for $\ell\geq 557$, we have $0.92\cdot\ell\leq\log(\ell\#)$. Checking numerically that for $1\leq\ell\leq 557$ we have $2^{\ell/2}\leq\ell\#$ we conclude that this inequality holds for all $\ell\geq 1$. Moreover we have the estimate $\sum_{p\leq\ell}\frac{\log(p)}{p-1}\leq 2\sum_{p\leq\ell}\frac{\log(p)}{p}\leq 2\log(\ell)$ following [RS62, Eq. (2.5), (2.11)]. Taking now $\ell$ such that $\ell\#\leq\Delta\leq(\ell+1)\#$ we obtain that $\delta(\Delta)\leq\frac{4}{\log(2)^{2}}\log(\Delta)^{2}=O(\log(\Delta)^{2})$. ∎
Proposition 5.15.
Given $u(x)\coloneqq\frac{a(x)}{b(x)}\in\mathbb{Q}(x)$ of degree $n$ and multiplicative height $H$, deciding if all solutions of $y^{\prime}(x)=u(x)y(x)$ are algebraic can be done using Algorithm 3 by performing $\tilde{O}(\Delta^{6}B)=\tilde{O}(H^{12n-6}n^{12n}3^{-3n})$ bit operations, where $\tilde{O}$ hides factors polynomial in $n$ and logarithmic in $H$.
Доказательство.
We compute $R(w),\Delta,$ and $\delta$ as described in the algorithm, and the maximal modulus of a root of $R(w)$, up to a constant factor by for instance setting $\tau=\frac{1}{2}$ in Proposition 5.11. This we will use as an upper bound on $B$. With this we obtain an upper bound on $\sigma$. Clearly the complexities of these computations are negligible compared to the computation of sufficiently many $p$-curvatures as described below.
By Theorem 1.3, it suffices to compute the $p$-curvatures up to $\sigma$, defined in its statement. Thus, we want to set $S=\sigma$ in Proposition 5.12. In accordance with Lemma 5.14 we have $\sigma=\tilde{O}(B\Delta^{6})$ where we omit logarithmic factors in $\Delta$. Cauchy’s bound on the roots of $R(w)=r_{n}w^{n}+\dots+r_{0}$ states that we can take $B$ at most $1+\max_{i}\left|\frac{r_{i}}{r_{n}}\right|$, hence $\sigma=\tilde{O}(\Delta^{5}H_{R})$, where $H_{R}$ denotes the multiplicative height of $R(w)$. Using estimates from Proposition 5.3 to estimate this height and $\Delta$ we find both $\Delta$ and $H_{R}$ to be in $\tilde{O}(H^{2n-1}n^{2n-1/2}3^{-n/2})$. Taking the $6$-th power of this quantity and combining with Proposition 5.12 concludes. ∎
We warn the reader that the multiplicative height $H$ appearing is exponential in the height $h$ appearing in other complexity estimates.
This finishes the proof of the complexity estimates in Theorem 1.3. Note that Algorithm 3 delays the computation of $B$ and $\sigma$, to after checking some amount of $p$-curvatures. This does not affect the complexity estimates presented here, but allows the algorithm to finish quickly for many examples, in which it returns transcendental. By Proposition 5.12, if, for example, we assume that among the first primes up to $C\cdot n^{k}\cdot h$ for some fixed $C>0$ and $k\in\mathbb{N}$ one $p$-curvature does not vanish, then the algorithm returns ‘‘Transcendental’’ in time $\tilde{O}(n^{k+2}h)$, where $h=\log_{2}(H)$. In practice we observe that for random examples, such an assumption can easily be made.
Our algorithm is to be compared with the Algorithm 1 based on finding rational roots of a polynomial, whose complexity is polynomial in the degree and linear in the height as stated in Corollary 5.10. This is vastly smaller than our exponential bounds in the degree $n$ and height $h$ to verify algebraicity. However, we expect rational functions yielding transcendental solutions to have $p$-curvatures that do not vanish for very small primes already, allowing Algorithm 3 to conclude much faster than Algorithm 1. Making an assumption as described above on the maximal expected first non-vanishing $p$-curvatures gives a heuristic about the differences in complexity in practice. This general behavior is highlighted on examples in the following section.
Remark 5.16.
As highlighted already in Remark 5.7, our algorithm extends to the case of $u(x)\in\overline{\mathbb{Q}}(x)\setminus\mathbb{Q}(x)$, however the complexity estimate of Proposition 5.15 does not hold because computations are more costly in number fields.
6 Implementation
6.1 Algorithmic Improvements
The following paragraphs discuss our implementation of Algorithm 3 in SageMath, that is available on GitHub with examples: https://github.com/plucas0/Honda.git.
The crucial algorithmic idea of the design of Algorithm 3 is to delay the moment when costly computations are done, that is the computation of the Rothstein-Trager resultant $R(w)$, when the degree of the input increases. By checking the ‘‘first’’ $p$-curvatures, in our implementation for primes $p$ up to $\Delta$, we try to return ‘‘Transcendental’’ before having to compute $R(w)$. In case no $p$-curvature vanishes until this point, the computation of further $p$-curvatures, potentially up to $\sigma$, is inevitable and this has much higher complexity than the computation of $R(w)$.
When computing the bound $B$ on the modulus of the roots of $R(w)$, the precision $\tau$ of the computation influences the complexity, see Proposition 5.11. In our case, computing all $p$-curvatures up to $\sigma$ already exceeds the capabilities of modern computers in most cases. For theoretical estimates, having $B$ multiplied by a constant factor does not change the complexity of the computations. Our implementation relies on native SageMath functions numerically computing zeroes of a polynomial, up to a fixed precision of $2^{-53}$.
Finally, the theoretical complexity of reducing polynomials modulo many primes $p$ simultaneously can be improved by adapting algorithms of polynomial multi-point evaluation for integers. However this is not done in our implementation, and reduction mod $p$ does not have a significant impact on computation time compared to the other operations performed by the algorithm.
Remark 6.1.
The delaying of the computation of the full Rothstein-Trager resultant is not as easily possible if $u(x)\in\overline{\mathbb{Q}}(x)\setminus\mathbb{Q}(x)$ as explained in Remark 5.1, and we do not know a method to delay this computation. For this reason, our implementation only treats the case of $u(x)\in\mathbb{Q}(x)$.
6.2 Competing Algorithms
6.2.1 Computation of $p$-curvatures
First of all we would like to ensure and to convince the reader that our implementation of Algorithm 4 of [BS09] to compute $p$-curvatures of first-order differential equations is indeed faster than Pagиs’ implementation of his algorithm [Pag21] applied to order one equations. The following timings in average were noted for both algorithms when asked to compute $p$-curvatures for primes up to $S$ on random polynomial inputs $a(x),b(x)\in\mathbb{Z}[x]$ of given multiplicative height $H$ and degrees n1n-1, respectively $n$.
Degree Height $S$ BS09 Pag21
33 22 $100$ 2020 ms $126$ ms
33 $2^{10}$ $100$ $23$ ms $120$ ms
33 $2^{80}$ $100$ $25$ ms $140$ ms
2020 22 $100$ $0.08$ s $13$ s
33 22 $1000$ $0.2$ s $1.2$ s
Таблица 1: Computation time of two algorithms computing $p$-curvatures.
We see that for ‘‘small’’ inputs Algorithm 4 is faster than Pagиs’ algorithm, and this remains the case when increasing any parameter.
6.2.2 Maple’s istranscendental
The maple package gfun [SZ94] has a command called istranscendental based on the algorithms described in [BSS25], aiming to prove transcendence of a given solution of a given differential equation. Among other things, it checks whether the singularities of the minimal differential operator annihilating a given function are all regular, and if its local exponents are rational. This is a necessary criterion for the algebraicity of the solution, so in case one of these conditions is violated, the command outputs true. For order one differential equations however, Proposition 3.3 ensures that the conditions are also sufficient for algebraicity. While istranscendental outputs FAIL in this case, we can actually conclude algebraicity and thus we are (ab)using the command as an implementation of Algorithm 2.
6.2.3 Rational Roots of the Rothstein-Trager resultant
In Section 5.2 we investigated the theoretical complexity of Algorithm 1 which decides algebraicity of the solutions of (2) by checking whether the Rothstein-Trager resultant of the coefficient $u(x)$ completely factors over $\mathbb{Q}$. This is equivalent to checking that the number of its rational roots is equal to its degree, as the algorithm describes. For this purpose, SageMath’s native roots command performs a full factorization in $\mathbb{Q}[x]$, while Maple has more efficient methods implemented. We will see in the timings in the next section, that the computation of the resultant is the computationally more complex task in practice.
6.3 Examples
6.3.1 Small Inputs
Tests of our implementation of Algorithm 3 on very small examples, with small inputs and algebraic output, quickly reach the limits of modern computers’ capabilities, as displayed in Table 2. We also compare our algorithm ($p$-curv) with istranscendental (ist) and the computation of the Rothstein-Trager resultant and its rational roots (RR) in Maple, and with computation of the resultant and performing its full factorization (fact) in SageMath.
$\cfrac{a(x)}{b(x)}$ $\sigma$ Output $p$-curv ist fact $R$
$\cfrac{3x-4}{2x^{2}-6x+4}$ $265$ A $120$ ms $45$ ms <1<1 ms $25$ ms
$\cfrac{7x^{2}-3x-4}{2x^{3}+4x^{2}-6x+4}$ $\approx 6\cdot 10^{27}$ T 55 ms $38$ ms <1<1 ms $30$ ms
$\cfrac{2x+1}{x^{2}+x+1}$ $1926284$ A 88min 99s 1919 ms <1<1 ms $24$ ms
$\cfrac{1}{x^{2}-4}$ $\approx 10^{11}$ A DNF $15$ ms <1<1 ms $22$ ms
Таблица 2: Output and computation time of our implementation of Algorithm 3 and timings of competing algorithms on a few examples. Here the output A stands for ‘‘Algebraic’’, and T stands for ‘‘Transcendental.’’
In particular, on the last line, it takes approximately $1$ minute to compute $p$-curvatures for primes $p$ up to $2\cdot 10^{5}$, an extrapolation indicates it would take more than $2$ years to compute them all up to $\sigma\approx 10^{11}$. At the same time it is obvious that the polynomial $x^{2}-4$ splits in $\mathbb{Q}[x]$, hence we would like to return ‘‘Algebraic’’ instantly.
6.3.2 Large Random Inputs
Certifying algebraicity using our approach is difficult. Both the theoretical complexity and the timings of our algorithm suggest so, already for small degree and height of the input. Proving transcendence, however, is in general much easier as we expect that in this case – except for very specific polynomials – the $p$-curvatures for some small primes $p$ will not vanish. This might seem counterintuitive when compared to numbers – for which proving transcendence is considered hard – but this behavior was already observed in the functional context in [BSS25]. By ‘‘small’’, we mean that we expect that it suffices to check a number of $p$-curvatures significantly smaller than the bound $\sigma$ of Theorem 1.3, but the latter is, to our knowledge, the best proven bound. Thus, in practice, our algorithm should be viewed rather only as a semi-algorithm in the sense that if the output is algebraic the algorithm will not terminate in reasonable time, except in very specific examples. The observed behavior worsens as the input degree and height increase, hence let us focus on the case where the output is transcendental.
It is expected that the solution of a first order differential equation with coefficient $u(x)=a(x)/b(x)$ for integer polynomials $a(x),b(x)$ of fixed degree $n\in\mathbb{N}_{\geq 1}$ with random integer coefficients in [H,H][-H,H], for $H\in\mathbb{N}_{>0}$ is transcendental, and checking that is possible by computation of very few $p$-curvatures. In Table 3, computation times of all algorithms mentioned in 6.2, including Algorithm 3, are compared on inputs consisting of two such polynomials, with the first polynomial’s degree strictly smaller than the second’s. The algorithms we compare are our implementation of Algorithm 3 in SageMath ($p$-curv), Maple’s istranscendental (ist), computing the Rothstein-Trager resultant (RT) and using Maple’s roots command for finding the rational roots of this resultant (RR), and computing the Rothstein-Trager resultant in SageMath and factoring it (fact). Of course the last algorithm is expected not to be competitive, still we display the timings to have a comparison in SageMath, the same system as our implementation.
RT+RR (Maple)
Degree Height $p$-curv ist RT RT+$R$ fact (Sage)
$10$ $2^{10}$ 11 ms $12$ ms 33 ms 33 ms <1<1 ms
2020 $^{10} 22$ ms $24$ ms 99 ms $10$ ms 44 ms
2020 $^{20} 22$ ms $25$ ms 1919 ms 2121 ms 77 ms
$40$ $2^{10}$ 44 ms $71$ ms $46$ ms $49$ ms $79$ ms
$40$ $2^{20}$ 55 ms $76$ ms $100$ ms $107$ ms $171$ ms
$80$ $2^{10}$ $0.1$ s $0.3$ s $0.3$ s $0.3$ s $2.4$ s
$80$ $2^{20}$ $0.1$ s $0.3$ s $0.6$ s $0.6$ s $5.0$ s
$160$ $2^{10}$ $0.4$ s $1.8$ s $2.4$ s $2.4$ s $83$ s
$160$ $2^{20}$ $0.4$ s $1.9$ s $3.9$ s $4.0$ s $182$ s
Таблица 3: Average computation time of various algorithms on random rational function inputs of prescribed degree and height.
The timings are given for random inputs $a(x),b(x)\in\mathbb{Z}[x]$ of multiplicative height at most $H$ and degree n1n-1, respectively, $n$, where Algorithm 3 returns ‘‘Transcendental.’’ The tests for degrees up to $40$ were performed on samples of hundreds of random rational functions, the size of the sample for the tests in degree $80$ and $160$ were respectively $100$ and $10$. In all our experiments, we observed that most examples return transcendental after computing a non-vanishing $p$-curvature with $p$ no greater than $17$. In such cases, we avoid the computation of the Rothstein-Trager polynomial and only compute its leading coefficients, the discriminant of $b(x)$, whereas the factorization approach cannot take this shortcut.
The timings of computing Rothstein-Trager resultants separately were only performed in Maple, which outperforms SageMath on this task. We observe that the time for finding all rational roots of the resultant is negligible compared to computing the resultant in the first place, and that our implementation outperforms any algorithm requiring the computation of the resultant. The timings nicely illustrate, that for generic examples our algorithm returns ‘‘Transcendental’’ quickly, as claimed at the end of Section 5.5.
One can hand-craft examples on which we need to test relatively large primes $p$ – at least compared to the generic behavior. In the case of a quadratic polynomial, this smallest prime for which it does not factor into a product of two linear factors is the smallest quadratic non-residue in its splitting field. Examples of quadratic fields with prescribed least quadratic non-residue can be found in [MT21]. From this one can construct coefficients $u(x)$ of differential equations, for which the first prime for which the $p$-curvature does not vanish is comparatively large. For instance, if $a(x)=1$, $b(x)=x^{2}-3818929$, the prime $p=2$ divides $\Delta=\operatorname{res}(b(x),b^{\prime}(x))$, and for all other primes up to $43$ the $p$-curvatures vanish, but not the $47$-curvature. Our algorithm runs in $80$ milliseconds on this example.
More generally, we already discussed the connection of our problem with bounding the least prime that does not split in a number field in Section 3.
A systematic construction of polynomials $R(w)$ with ‘‘large’’ smallest prime $p$ for which the reduction of $R(w)$ does not split completely, or of coefficients $u(x)$ of differential equations for which the first non-vanishing $p$-curvature is high, is not known to us.
Список литературы
  • [AK19] Jeoung-Hwan Ahn and Soun-Hi Kwon ‘‘An Explicit Upper Bound for the Least Prime Ideal in the Chebotarev Density Theorem’’ In Annales de l’Institut Fourier 69.3, 2019, pp. 1411–1458 DOI: 10.5802/aif.3274
  • [Ax71] James Ax ‘‘On Schanuel’s conjectures’’ In Ann. of Math. (2) 93, 1971, pp. 252–268 DOI: 10.2307/1970774
  • [BCR24] Alin Bostan, Xavier Caruso and Julien Roques ‘‘Algebraic Solutions of Linear Differential Equations: An Arithmetic Approach’’ In Bulletin of the American Mathematical Society 61.4, 2024, pp. 609–658 DOI: 10.1090/bull/1835
  • [BCS15] Alin Bostan, Xavier Caruso and Йric Schost ‘‘A fast algorithm for computing the $p$-curvature’’ In ISSAC’15—Proceedings of the 2015 ACM International Symposium on Symbolic and Algebraic Computation ACM, New York, 2015, pp. 69–76
  • [BCS16] Alin Bostan, Xavier Caruso and Йric Schost ‘‘Computation of the similarity class of the $p$-curvature’’ In Proceedings of the 2016 ACM International Symposium on Symbolic and Algebraic Computation ACM, New York, 2016, pp. 111–118 DOI: 10.1145/2930889.2930897
  • [BH89] F. Beukers and G. Heckman ‘‘Monodromy for the Hypergeometric Function $_{n}F_{n-1}$’’ In Inventiones Mathematicae 95.2, 1989, pp. 325–354 DOI: 10.1007/BF01393900
  • [Bos+17] Alin Bostan et al. ‘‘Algorithmes Efficaces en Calcul Formel’’ 686 pages. Imprimй par CreateSpace. Aussi disponible en version йlectronique Palaiseau: Frйdйric Chyzak (auto-йdit.), 2017 URL: https://hal.archives-ouvertes.fr/AECF/
  • [Bou97] Auguste Boulanger ‘‘Contribution а l’йtude des йquations diffйrentielles linйaires et homogиnes intйgrables algйbriquement’’, 1897
  • [BS09] Alin Bostan and Йric Schost ‘‘Fast algorithms for differential equations in positive characteristic’’ In Proceedings of the 2009 international symposium on symbolic and algebraic computation, ISSAC 2009, Seoul, July 28–31, 2009. New York, NY: Association for Computing Machinery (ACM), 2009, pp. 47–54 DOI: 10.1145/1576702.1576712
  • [BSS25] Alin Bostan, Bruno Salvy and Michael F. Singer ‘‘On Deciding Transcendence of Power Series’’, 2025 DOI: 10.48550/arXiv.2504.16697
  • [CC85] D.. Chudnovsky and G.. Chudnovsky ‘‘Applications of Padй approximations to the Grothendieck conjecture on linear differential equations’’ In Number theory (New York, 1983–84) 1135, Lecture Notes in Math. Springer, Berlin, 1985, pp. 52–100 DOI: 10.1007/BFb0074601
  • [Chu83] G.. Chudnovsky ‘‘On the method of Thue-Siegel’’ In Ann. Math. (2) 117, 1983, pp. 325–382 DOI: 10.2307/2007080
  • [Fid85] Charles M. Fiduccia ‘‘An efficient formula for linear recurrences’’ In SIAM J. Comput. 14.1, 1985, pp. 106–112 DOI: 10.1137/0214007
  • [FY24] Florian Fьrnsinn and Sergey Yurkevich ‘‘Algebraicity of Hypergeometric Functions with Arbitrary Parameters’’ In Bulletin of the London Mathematical Society, 2024, pp. 23 pp. DOI: 10.1112/blms.13103
  • [Her73] Ch. Hermite ‘‘Sur la fonction exponentielle’’ In C. R. Acad. Sci., Paris 77, 1873, pp. 18–24
  • [Her93] Ch. Hermite ‘‘Sur la gйnйralisation des fractions continues algйbriques, Extrait d’une lettre а M. Pincherlй’’ In Annali di Mat. (2) 21, 1893, pp. 289–308 DOI: 10.1007/BF02420446
  • [Hon81] Taira Honda ‘‘Algebraic Differential Equations’’ In Symposia Mathematica XXIV (Sympos., INDAM, Rome, 1979) London-New York: Academic Press, 1981, pp. 169–204
  • [Jac37] Nathan Jacobson ‘‘Abstract Derivation and Lie Algebras’’ In Transactions of the American Mathematical Society 42.2, 1937, pp. 206 DOI: 10.2307/1989656
[…]
  • [Jag64a] H. Jager ‘‘A multidimensional generalization of the Padй table. IV, V, VI’’ In Nederl. Akad. Wet., Proc., Ser. A 67, 1964, pp. 227–249
  • [Kat72] Nicholas M. Katz ‘‘Algebraic Solutions of Differential Equations ($p$-Curvature and the Hodge Filtration)’’ In Inventiones Mathematicae 18.1-2, 1972, pp. 1–118 DOI: 10.1007/BF01389714
  • [Kat82] Nicholas M. Katz ‘‘A Conjecture in the Arithmetic Theory of Differential Equations’’ In Bulletin de la Sociйtй mathйmatique de France 79, 1982, pp. 203–239 DOI: 10.24033/bsmf.1960
  • [Kro80] Leopold Kronecker ‘‘Ьber die Irreducibilitдt von Gleichungen’’ In Monatsberichte der Kцniglichen Preussische Akademie des Wissenschaften zu Berlin, 1880, pp. 155–163
  • [KW22] Habiba Kadiri and Peng-Jie Wong ‘‘Primes in the Chebotarev Density Theorem for All Number Fields (with an Appendix by Andrew Fiori)’’ In Journal of Number Theory 241, 2022, pp. 700–737 DOI: 10.1016/j.jnt.2022.03.012
  • [Li12] Xiannan Li ‘‘The Smallest Prime That Does Not Split Completely in a Number Field’’ In Algebra & Number Theory 6.6, 2012, pp. 1061–1096 DOI: 10.2140/ant.2012.6.1061
  • [LMO79] J.. Lagarias, H.. Montgomery and A.. Odlyzko ‘‘A Bound for the Least Prime Ideal in the Chebotarev Density Theorem’’ In Inventiones Mathematicae 54.3, 1979, pp. 271–296 DOI: 10.1007/BF01390234
  • [Loo83] Rьdiger Loos ‘‘Computing rational zeros of integral polynomials by $p$-adic expansion’’ In SIAM J. Comput. 12.2, 1983, pp. 286–293 DOI: 10.1137/0212017
  • [Mah68] K. Mahler ‘‘Perfect systems’’ In Compos. Math. 19, 1968, pp. 95–166 URL: https://eudml.org/doc/88959
  • [MT21] Kevin McGown and Enrique Treviсo ‘‘The least quadratic non-residue’’ In Mexican mathematicians in the world—trends and recent contributions 775, Contemp. Math. Amer. Math. Soc., 2021, pp. 205–231 DOI: 10.1090/conm/775/15593
  • [Mur94] V. Murty ‘‘The Least Prime Which Does Not Split Completely’’ In Forum Mathematicum 6.6, 1994 DOI: 10.1515/form.1994.6.555
  • [Pad01] H. Padй ‘‘Sur l’expression gйnйrale de la fraction rationnelle approchйe de $(1+x)^{m}$.’’ In C. R. Acad. Sci., Paris 132, 1901, pp. 754–756
  • [Pad92] H. Padй ‘‘Sur la reprйsentation approchйe d’une fonction par des fractions rationnelles’’ In Ann. Sci. Йc. Norm. Supйr. (3) 9, 1892, pp. 3–93 DOI: 10.24033/asens.378
  • [Pad94] H. Padй ‘‘Sur la gйnйralisation des fractions continues algйbriques’’ In C. R. Acad. Sci., Paris 118, 1894, pp. 848–850
  • [Pag21] Raphaлl Pagиs ‘‘Computing Characteristic Polynomials of $p$-Curvatures in Average Polynomial Time’’ In Proceedings of the 2021 International Symposium on Symbolic and Algebraic Computation Virtual Event Russian Federation: ACM, 2021, pp. 329–336 DOI: 10.1145/3452143.3465524
  • [Pai87] Paul Painlevй ‘‘Sur es йquations diffйrentielles linйaires’’ In Comptes Rendus de l’Acadйmie des Sciences, Paris 105, 1887, pp. 1829–1832
  • [Put96] Marius Put ‘‘Reduction modulo $p$ of differential equations’’ In Indagationes Mathematicae 7.3, 1996, pp. 367–387 DOI: https://doi.org/10.1016/0019-3577(96)83726-8
  • [Ris70] Robert H. Risch ‘‘The Solution of the Problem of Integration in Finite Terms’’ In Bulletin of the American Mathematical Society 76.3, 1970, pp. 605–608 DOI: 10.1090/S0002-9904-1970-12454-5
  • [Rob55] Herbert Robbins ‘‘A Remark on Stirling’s Formula’’ In The American Mathematical Monthly 62.1 [Taylor Francis, Ltd., Mathematical Association of America], 1955, pp. 26–29 URL: http://www.jstor.org/stable/2308012
  • [Rot76] Michael Rothstein ‘‘Aspects of Symbolic Integration and Simplification of Exponential and Primitive Functions’’, 1976
  • [Rot77] Michael Rothstein ‘‘A New Algorithm for the Integration of Exponential and Logarithmic Functions’’ In Proceedings of the 1977 MACSYMA Users’ Conference, 1977
  • [RS62] J. Rosser and Lowell Schoenfeld ‘‘Approximate formulas for some functions of prime numbers’’ In Illinois Journal of Mathematics 6.1 Duke University Press, 1962, pp. 64–94 DOI: 10.1215/ijm/1255631807
  • [Sch82] Arnold Schцnhage ‘‘The fundamental theorem of algebra in terms of computational complexity’’ In Manuscript. Univ. of Tьbingen, Germany, 1982
  • [Sin79] Michael F. Singer ‘‘Algebraic Solutions of $n$th Order Linear Differential Equations’’ In Proceedings of the Queen’s University 1979 Conference on Number Theory 54, Queens Papers in Pure and Applied Mathematics, 1979, pp. 379–420
  • [Sta80] R.P. Stanley ‘‘Differentiably Finite Power Series’’ In European Journal of Combinatorics 1.2, 1980, pp. 175–188 DOI: 10.1016/S0195-6698(80)80051-5
  • [SZ94] Bruno Salvy and Paul Zimmermann ‘‘GFUN: A Maple Package for the Manipulation of Generating and Holonomic Functions in One Variable’’ In ACM Transactions on Mathematical Software 20.2, 1994, pp. 163–177 DOI: 10.1145/178365.178368
  • [Tra76] Barry M. Trager ‘‘Algebraic Factoring and Rational Function Integration’’ In Proceedings of the Third ACM Symposium on Symbolic and Algebraic Computation - SYMSAC ’76 Yorktown Heights, New York, United States: ACM Press, 1976, pp. 219–226 DOI: 10.1145/800205.806338
  • [Tsc26] N. Tschebotareff ‘‘Die Bestimmung der Dichtigkeit einer Menge von Primzahlen, welche zu einer gegebenen Substitutionsklasse gehцren’’ In Mathematische Annalen 95.1 Springer Science and Business Media LLC, 1926, pp. 191–228 DOI: 10.1007/bf01206606
  • [VV00] Jeffrey D Vaaler and Josй Felipe Voloch ‘‘The Least Nonsplit Prime in Galois Extensions of $\mathbb{Q}$’’ In Journal of Number Theory 85.2, 2000, pp. 320–335 DOI: 10.1006/jnth.2000.2551
  • [vzGG13] Joachim Gathen and Jьrgen Gerhard ‘‘Modern Computer Algebra’’ Cambridge: Cambridge university press, 2013
  • [Чеб23] Н Чеботарев ‘‘Определение плотности совокупности простых чисел, принадлежащих к заданному классу подстановок. I’’ In Известия Российской Академии Наук 17.1-18, VI, 1923, pp. 205–230
  • [Чеб23a] Н Чеботарев ‘‘Определение плотности совокупности простых чисел, принадлежащих к заданному классу подстановок. II’’ In Известия Российской Академии Наук 17.1-18, VI, 1923, pp. 231–250
Faculty of Mathematics, University of Vienna, Oskar-Morgenstern-Platz 1, 1090, Vienna, Austria
CNRS; Laboratoire de Mathйmatiques de Versailles, UMR-8100, UVSQ, 45 avenue des Йtats-Unis 78000, Versailles CEDEX, France